### v3

# New Bounds on Quotient Polynomials with Applications to Exact Divisibility and Divisibility Testing of Sparse Polynomials

Abstract
We prove that for polynomials $f,g\in\mathbb{C}[x]$ such that $g$ divides $f$, the $\ell_{2}$-norm of the quotient polynomial $f/g$ is bounded by $\left\|{f/g}\right\|_{2}\leq\left\|{f}\right\|_{1}\cdot\tilde{O}(\left\|{g}\right\|_{0}^{3}\deg^{2}{f})^{\left\|{g}\right\|_{0}-1}$, where $\|g\|_{0}$ is the sparsity of $g$ (the number of its non-zero coefficients). This improves upon the previously known exponential (in $\deg(f)$) bounds for general polynomials. Our result implies that the trivial long division algorithm runs in quasi-linear time relative to the input size and number of terms of the quotient polynomial $f/g$, thus solving a long-standing problem on exact divisibility of sparse polynomials. Prior to our work it was not even known that the representation size of the quotient polynomial is sub-quadratic in its number of terms, i.e., that it is sub-quadratic in $\deg{f}$.
We also study the problem of bounding the number of terms of $f/g$ in some special cases. When $f,g\in\mathbb{Z}[x]$ and $g$ is a cyclotomic-free (i.e., it has no cyclotomic factors) trinomial, we prove that $\left\|{f/g}\right\|_{0}\leq O(\left\|{f}\right\|_{0}\text{size}\left(f\right)^{2}\cdot\log^{6}{\deg{g}})$. When $g$ is a cyclotomic-free binomial, we prove that the sparsity is at most $O(\left\|{f}\right\|_{0}(\log{\left\|{f}\right\|_{0}}+\log{\left\|{f}\right\|_{\infty}}))$. Both upper bounds are polynomial in the input-size. Without assuming cyclotomic-freeness, there are simple examples that show that both sparsities can be exponentially large. Leveraging these results, we provide a polynomial-time algorithm for deciding whether a cyclotomic-free trinomial divides a sparse polynomial over the integers. In addition, we show that testing divisibility by a binomial can be done in $\tilde{O}(n^{2})$ time over $\mathbb{Z}$ and over finite fields.
As our last result, we present a polynomial time algorithm for testing divisibility by pentanomials over small finite fields when $\deg{f}=\tilde{O}(\deg{g})$.
All our algorithms are deterministic.
1 Introduction
The sparse representation of polynomials is widely used in computer algebra as it is the most natural and succinct representation when polynomials are far from being “dense”, i.e., when most of their coefficients are zero, or, equivalently, when they are supported on a small number of monomials [vzGG13, DST93]. In this representation, a polynomial is described by a list of coefficient and degree pairs that correspond to its terms with non-zero coefficients. For example, a polynomial $f\in R[x]$ over the ring $R$ with $t$ non-zero coefficients, $f=\sum_{i=1}^{t}{c_{i}x^{e_{i}}}$, is represented as $\left\{(c_{i},e_{i})\mid i=1\ldots t\right\}$. Particularly interesting cases are when $R=\mathbb{Z}$ or $R$ is a finite field. When working over the integers, the size of the sparse representation of a polynomial $f\in\mathbb{Z}[x]$, denoted $\text{size}\left(f\right)$, is
$\text{size}\left(f\right)=\left\|{f}\right\|_{0}(\log_{2}{\left\|{f}\right\|_{\infty}}+\log_{2}{\deg{f}}),$
where $\left\|{f}\right\|_{0}$ and $\left\|{f}\right\|_{\infty}$ are the number of non-zero coefficients in $f$, and the maximal coefficient of $f$ (in absolute value), respectively.111 As we think of $\left\|{f}\right\|_{0}$ as being considerably smaller than $\log_{2}{\left\|{f}\right\|_{\infty}}$ and $\log_{2}{\deg{f}}$, we do not pay attention to the case where bit lengths of coefficients vary. Similarly, over a finite field $\mathbb{F}$ the representation size is
$\text{size}\left(f\right)=\left\|{f}\right\|_{0}(\log_{2}{\left|\mathbb{F}\right|}+\log_{2}{\deg{f}}).$
This makes the sparse representation particularly relevant to applications in which the polynomial has a large degree and most of its coefficients are zero. In particular, it is used by computer algebra systems and libraries such as Maple [MP13], Mathematica [Wol99], Sage [S+08], and Singular [Sch03].
In the sparse representation, the degree could be exponentially larger than the representation size. Algorithms that run in polynomial time in terms of the sparse-representation size (i.e. number of terms, their bit complexity and the logarithm of the degree) are called super-sparse or lacunary polynomial algorithms.
[…]
One of the long-standing open problems concerns exact divisibility of sparse polynomials. There are two versions to this problem. The decision problem asks, given two sparse polynomials $f$ and $g$, to decide whether $g$ divides $f$. The search problem, or the exact division problem, asks to compute the quotient (also called co-factor) polynomial $f/g$ when it is known that $g$ divides $f$. The surveys of Davenport and Carette [DC09] and of Roche [Roc18] thoroughly discuss these problems.
The two problems are closely related but also differ in what we can hope to achieve. In the decision problem, the output is a single bit, allowing us to analyze the complexity in terms of the input size. Conversely, in the search problem, the output size depends on the number of monomials in $f/g$, which may require an exponential number of bits to represent. Next, we will discuss each of these problems separately.
1.1 Exact Division
There are two difficulties in the exact division problem. The first stems from the fact that $f/g$ can have exponentially (in the input-size) many monomials. The second concerns the height (i.e. the maximum among the absolute values of the coefficients) of the quotient polynomial, when working over the integers. Prior to this work the only bounds on the size of its coefficients were exponential in $\deg{f}$ (see [Mig12]). [^2] These results are tight when $g$ is not a sparse polynomial. In particular, Davenport and Carette note that even in the case where $g$ exactly divides $f$, we need to perform $\Omega(\|f/g\|_{0}\cdot\|g\|_{0})$ operations on coefficients, and the number of bits required to represent a coefficient may be as large as $\deg f$ [DC09, Section III.B]. Therefore, in terms of the input size and $\|f/g\|_{0}$ the only provable bounds on the running time of the exact division algorithm were of the form
$\Omega(\|f/g\|_{0}\cdot\|g\|_{0}\cdot\deg f).$ (1.1)
1.2 Divisibility Testing
Recall that in the divisibility testing problem the goal is to determine whether $g$ divides $f$ without having to compute the quotient and remainder. Thus, we seek a polynomial-time algorithm in the size of the input. It is easy to see that when $\deg{g}$ is small (or, similarly, $\deg{f}-\deg{g}$ is small), or when the quotient polynomial is sparse, the problem becomes easy.
However, in the general case, it is only known that, assuming the Extended Riemann Hypothesis (ERH), the divisibility problem is in coNP [GKO92]. This led Davenport and Carette to pose the following challenge:
Open Problem 1 (Challenge 3 of [DC09]).
Find a class of instances for which the question "does g divide f?" is NP-complete, or find an algorithm for the divisibility of polynomials which is polynomial-time. Failing this - find an algorithm for the divisibility of cyclotomic-free polynomials which is polynomial-time.
In fact, even the next problem, that was stated in [Roc18] following [DC09], is still open. [^3] As usual, $\tilde{O}(t)=O\left(t\cdot\log^{O(1)}t\right)$ for a running time function $t$, and when no basis is mentioned then the log is taken in the natural base $e$
Open Problem 2 ([Roc18, Open Problem 3]).
Given two sparse polynomials $f,g\in R[x]$, develop an algorithm which determines whether $g$ divides $f$ exactly, using $\tilde{O}(T\log{\deg{f}})$ ring and bit operations, where $T$ is the total number of terms of $f$, $g$, $(f\text{ quo }g)$ and $(f\text{ rem }g)$ (the quotient and remainder).
Note that 2 asks for an algorithm whose running time may be exponential in the input size.
There are several reasons to believe that divisibility testing is a hard problem. In [Pla84], Plaisted studied the problem of divisibility testing of sparse polynomials and proved that determining whether the quotient polynomial has a non-zero constant term is NP-hard and that determining the degree of the remainder polynomial is also NP-hard. Divisibility testing of sparse polynomials is a special case of a more general problem of divisibility testing of arithmetic circuits in one variable (of arbitrary degree). This problem too was shown to be NP-hard - as mentioend, testing whether $x^{n}-1$ divides a product of sparse polynomials [Pla77] is NP-Hard. While these results do not prove that the divisibility problem is NP-hard, they show that closely related problems are difficult to solve.
Davenport and Carette highlighted polynomials with cyclotomic factors in 1 because, since Plaisted’s work [Pla77], all NP-hard problems related to sparse polynomials involve cyclotomic polynomials. For example, many problems are known to be NP-hard when the polynomials involved can have cyclotomic factors, such as deciding whether there exists a non-trivial common divisor of two polynomials, testing whether a polynomial has a cyclotomic factor [Pla84], and determining if a polynomial is square-free [KS99]. Notably, it remains an open question whether these problems, and the NP-hard problem mentioned above, remain NP-hard when considering cyclotomic-free polynomials. For further discussion on the role of cyclotomic polynomials and the difficulties they pose, see the surveys [DC09, Roc18].
There are multiple obstacles hindering natural approaches, including long-division, computing the remainder, or testing evaluations, from efficiently solving the divisibility testing problem. We first note that the quotient of two sparse polynomials may not be sparse. For example, consider the equation
$\frac{x^{n}-1}{x-1}=\sum_{0\leq i<n}{x^{i}}.$
Here the size of the input is $O(\log n)$ whereas the size of the output is $\Omega(n)$. This rules out the simple long division algorithm and requires any solution to decide whether such a quotient exists without computing it explicitly. Similarly, we note that the remainder of the division of two sparse polynomials may not be sparse as well, and, moreover, may contain doubly exponential large coefficients. For example, consider
$x^{(n+1)n}\bmod x^{n+1}-x-1=\sum_{0\leq i\leq n}{\binom{n}{i}x^{i}}.$
Here the input size is $O(\log n)$, however, the coefficient of $x^{n/2}$ of the remainder is of size $2^{n-O(\log n)}$, which is doubly exponential in the input size. This rules out any approach that computes the remainder explicitly. Nevertheless, a solution to the divisibility testing problem has to decide whether the remainder is zero or not. Another challenge arises from the inefficiency of evaluating polynomials: the evaluation of a sparse polynomial $f\in\mathbb{Z}[x]$ on an algebraic number $a$ may have an exponential bit complexity, unless $\left|a\right|=1$ or $a=0$. Thus, natural approaches to the divisibility problem are not effective when considering sparse polynomials of high degree. Prior to our work, the best upper bounds on the size of the coefficients of the quotient polynomial, when $g$ exactly divides $f$, were exponentially large in the degree of $f$, see [Mah62], [BG06, Chapter 1.6] and [Mig12, Proposition 4.9]. Thus, it was previously unknown whether coefficients of the quotient polynomial have “reasonable” representation size, even when the divisor polynomial has a bounded number of terms and the remainder is zero.
Some of the obstacles described above disappear when working over finite fields, such as having too large coefficients, or not being able to evaluate efficiently. However, there are reasons to believe that divisibility testing is a hard problem over finite fields as well. Bi, Cheng, and Rojas proved that with respect to BPP-reductions, deciding whether two sparse polynomials have non-trivial common divisor or whether a sparse polynomial is square-free is NP-Hard. [BCR16]. Other related problems such as computing the number of common roots of $n$ sparse polynomials $f_{1},\ldots,f_{n}\in\mathbb{F}_{p}[x]$, of degrees $\exp(n)$, were shown to be #$\#P$ complete [Qui86, vzGKS93].
This work:
We make significant progress toward solving 2 in the exact case and improve the bound given in Equation (1.1) on the running time of the exact division algorithm.
Specifically, we prove that, when $g$ exactly divides $f$ it holds that
$\log\|f/g\|_{\infty}\leq\log\|f\|_{\infty}+\tilde{O}\left(\|g\|_{0}\right)\cdot\log\deg f.$
For example, our result shows that when $g$ has a bounded number of terms, the coefficients of the quotients are at most polynomial in $\deg f$, thus yielding an exponential improvement over classical results [Mah62], [BG06, Chapter 1.6] and [Mig74].
An implication of this result is that, in the exact case, the running time of the long division algorithm is
$\tilde{O}(\|f/g\|_{0}\cdot\|g\|_{0}\cdot(\log\|f\|_{\infty}+\log\|g\|_{\infty}+\|g\|_{0}\cdot\log\deg f)),$ (1.2)
where earlier results could only prove an upper bound of the form $\Omega(\|f/g\|_{0}\cdot\|g\|_{0}\cdot\deg f)$ (see Equation (1.1)) [DC09].
[…]
We next discuss more related work and then present our results.
1.3 Related Work
1.3.1 Bounds on the Coefficients of the Quotient Polynomial
Let $f,g,h\in\mathbb{F}[x]$ such that $f=gh$. As mentioned in [GGPdC21, Lemma 2.12], a simple induction on the long division algorithm yields the following bound on the coefficients of $h$:
$\left\|{h}\right\|_{\infty}\leq\left\|{f}\right\|_{\infty}(\left\|{g}\right\|_{\infty}+1)^{\frac{\left\|{h}\right\|_{0}+1}{2}}.$ (1.3)
This bounds the representation size of the coefficients of $h$ by a linear function of $\left\|{h}\right\|_{0}$. However, when the quotient $h$ is not sparse, this bound becomes exponential in the degree of $h$, which can be exponential in the representation size of $f$ and $g$.
The best bound for this problem is due to Mignotte [Mig74], which is notable for being used in the polynomial factorization algorithm of Lenstra, Lenstra, and Lovász [LLL82].
Theorem 1.4 (Mignotte’s Bound [Mig74, Theorem 2]).
Let $f,g,h\in\mathbb{Q}[x]$ such that $f=gh$. Then,
$\left\|{h}\right\|_{1}\leq 2^{\deg{f}}\frac{\left\|{f}\right\|_{2}}{\left\|{g}\right\|_{1}}.$
Since in our case $\left\|{g}\right\|_{1}$ is relatively small, it is also worth noting another version of this bound which states that
$\left\|{h}\right\|_{1}\leq 2^{\deg{h}}\left\|{f}\right\|_{2}$ (1.5)
for any factor $h$ of $f$. Discussions about this classical inequality can be found in [Mig88, Section 2] and [Mig12, Chapter 4]. These bounds are also proven to be somewhat tight; for any $d\in\mathbb{N}$ there exists $h\in\mathbb{Z}[x]$ of degree $d$ and $f\in\mathbb{Z}[x]$ such that
$\left\|{h}\right\|_{1}\geq\Omega\left(\frac{2^{d}}{d^{2}\log{d}}\right)\left\|{f}\right\|_{2}.$
Another classical bound is known as Gel’fond’s Lemma (see [Mah62] and [BG06, Lemma 1.6.11]).
Theorem 1.6 ( Gel’fond’s Lemma).
Let $f,g,h\in\mathbb{Q}[x]$ such that $f=gh$. Then,
$\left\|{h}\right\|_{\infty}\leq 2^{\deg{f}}\frac{\left\|{f}\right\|_{\infty}}{\left\|{g}\right\|_{\infty}}.$ (1.7)
However, none of the above bounds is useful for divisibility testing when the degrees of $f,g$ and $h$ are exponentially large.
1.3.2 Exact Division of Sparse Polynomials
Monagan and Pearce [MP11, MP07] gave an algorithm for computing the quotient and remainder that performs $O(\|f/g\|_{0}\cdot\|g\|_{0})$ many ring operations. This algorithm ignores the bit complexity of the representation as, by the examples given above, the coefficients of the remainder polynomial can be doubly exponentially large in the input size. Their algorithm runs well when the result has small representation size.
In [GGPdC21, GGPdCR22] Giorgi, Grenet, Perret du Cray and Roche studied the exact divisibility problem over both finite fields and integers. They proposed randomized algorithms that run in quasi-linear time in the input size and representation size of the quotient polynomial. However, prior to our work, over the integers (or large finite fields), the best bounds on the coefficient size of the quotient were exponential in the input size. Hence, the only provable bound on the running time, in terms of the input and output size, was known to be quadratic. Moreover, prior to our work no deterministic quasi-linear time algorithm in the representation size of the quotient polynomial was known.
[…]
1.3.3 Testing Divisibility of Sparse Polynomials
[…]
A notable related result is due to Filaseta and Schinzel [FS04], who gave a polynomial time algorithm for deciding whether a sparse polynomial over the integers, with a bounded number of non-zero terms (i.e. $O(1)$ many terms), has a cyclotomic factor. In the general case, when the sparsity is not bounded, this problem was shown to be NP-complete [CTV10].
In [FGS08] Filaseta, Granville, and Schinzel showed that if $f$ and $g$ have bounded sparsities, bounded heights, and at least one of them is cyclotomic-free, then testing whether $g$ divides $f$ can be solved in linear time by computing their greatest common divisor.
A recent work by Giorgi, Grenet, and Perret du Cray [GGPdC21] shows that there exists a polynomial time algorithm for divisibility testing of sparse polynomials, in several special cases. Roughly, they proved that when there is a certain gap between the highest and second highest exponents in the divisor polynomial $g$ then the sparsity of the quotient polynomial can be bounded (which enables them to check divisibility using long division). Another special case that they handle is when $g=g_{0}+x^{k}g_{1}$ where $k=\Omega(\deg{f}-\deg{g})$ and $\deg{g_{0}}$ (or $\deg{g_{1}}$) is poly-logarithmic in the input size. A special case of their result applies to testing divisibility by trinomials; specifically, it implies that there exists a polynomial time algorithm that decides whether $g\mid f$ when $g$ is a trinomial and $\deg{f}=O(\deg{g})$. We obtain such a result for all cyclotomic-free trinomials $g$ and sparse polynomials $f$, regardless of their degrees.
1.4 Our Results
Our first main result gives a new bound on the $\ell_{2}$-norm of a quotient of two sparse polynomials over $\mathbb{C}$, thus improving the classical bounds of Mignotte [Mig74] and Gel’fond [Mah62] on the height of the quotient polynomial when $g$ is sparse.
Theorem 1.8 (Restatement of Theorem 3.10).
Let $f,g,h\in\mathbb{C}[x]$ such that $f=gh$ and $g$ is monic. Then,
$\left\|{h}\right\|_{2}\leq\left\|{f}\right\|_{1}\tilde{O}\left(\left\|{g}\right\|_{0}^{3}\deg^{2}{f}\right)^{\left\|{g}\right\|_{0}-1}.$
As opposed to the bounds given in Equations (1.3), (1.5) and (1.7), our bound on the coefficients of $h$ is not exponential in the degree of $g$, but rather in its sparsity, which for sparse polynomials may be exponentially smaller. Even more surprisingly, the bound in (1.9) does not depend on the height of $g$, in spite of the “feedback process” in the long division algorithm. We also get the following bound as an immediate corollary of Theorem 1.8.
$\log_{2}{\left\|{h}\right\|_{\infty}}\leq\log_{2}{\left\|{f}\right\|_{\infty}}+\tilde{O}(\|g\|_{0})\cdot\log_{2}{\deg{f}}.$ (1.9)
Thus, when $\|g\|_{0}$ is small this gives an exponential improvement over the bounds of Mignotte and Gel’fond (Theorems 1.41.6). In addition, Theorem 1.8 gives an exponential separation between exact and non-exact division since $\log_{2}{\left\|{h}\right\|_{\infty}}$ can be as large as $\Omega(\deg{f}\cdot\log_{2}{\left\|{g}\right\|_{\infty}})$ when allowing non-zero remainder.
As a corollary of Theorem 1.8 we obtain the following guarantee on the performance of the long division algorithm in the exact case.
Theorem 1.10.
When the division is exact, the time complexity of the long division algorithm for polynomials $f,g\in\mathbb{Z}[x]$ is
$\tilde{O}\left(\left\|{f/g}\right\|_{0}\cdot\left\|{g}\right\|_{0}\cdot\left(\log{\left\|{g}\right\|_{\infty}}+\log{\left\|{f}\right\|_{\infty}}+\left\|{g}\right\|_{0}\log{\deg{f}}\right)\right).$
[…]
Combining the bound in Theorem 1.8 with Theorem 1.3 of [GGPdCR22] we obtain the following corollary.
Corollary 1.11.
There exists a randomized algorithm that, given two sparse polynomials $f,g\in\mathbb{Z}[x]$, such that $g$ divides $f$, the algorithm outputs the quotient polynomial $f/g$, with probability at least $2/3$. The running time of the algorithm is
$\tilde{O}\left((\left\|{f}\right\|_{0}+\left\|{g}\right\|_{0}+\left\|{f/g}\right\|_{0})\cdot\left(\log{\left\|{g}\right\|_{\infty}}+\log{\left\|{f}\right\|_{\infty}}+\left\|{g}\right\|_{0}\log{\deg{f}}\right)\right).$
Next, we prove new bounds on the sparsity of the quotient of two polynomials over $\mathbb{Z}$ when $g$ is a cyclotomic-free binomial or a cyclotomic-free trinomial.
Theorem 1.12 (Restatement of Theorem 4.6).
Let $f,g,h\in\mathbb{Z}[x]$ such that $f=gh$. If $g$ is cyclotomic-free, and contains two terms then the sparsity of $h$ is at most
$O\left(\left\|{f}\right\|_{0}(\log{\left\|{f}\right\|_{0}}+\log{\left\|{f}\right\|_{\infty}})\right).$
Theorem 1.13 (Restatement of Theorem 4.17).
Let $f,g,h\in\mathbb{Z}[x]$ such that $f=gh$. If $g$ is cyclotomic-free and contains three terms, then the sparsity of $h$ is at most
$O\left(\left\|{f}\right\|_{0}\cdot\text{size}\left(f\right)^{2}\cdot\log^{6}{\deg{g}}\right).$
The above bounds give rise to explicit algorithms for testing the divisibility of polynomials over the integers by a binomial and by a cyclotomic-free trinomial. In what follows, $n$ will denote the size of the input, i.e., $\text{size}\left(f\right)+\text{size}\left(g\right)$.
Theorem 1.14.
Let $f,g\in\mathbb{Z}[x]$ be two polynomials given in the sparse representation, such that $g$ contains two terms. Then, there exists an algorithm that decides whether $g\mid f$ (and if so, computes the quotient $f/g$) in $\tilde{O}(n^{2})$ time.
Theorem 1.15.
Let $f,g\in\mathbb{Z}[x]$ be two polynomials given in the sparse representation, such that $g$ is cyclotomic-free and contains three terms. Then, there exists an algorithm that decides whether $g\mid f$ (and if so, computes the quotient $f/g$) in $\tilde{O}(n^{10})$ time.
Theorem 1.15 gives a positive answer to 1 for the case of trinomials over the integers.
Finally, we address the problem of polynomial divisibility over finite fields and show that testing divisibility by a binomial can be solved efficiently.
Theorem 1.16.
Let $f,g\in\mathbb{F}_{p}[x]$ be two polynomials given in the sparse representation, such that $g$ contains two terms. Then, there exists an algorithm that decides whether $g\mid f$ in $\tilde{O}(n^{2})$ time.
Binomials are a special case of a large class of polynomials, for which our algorithm also applies.
Remark 1.17.
[…]
As our last result, we present a new algorithm for divisibility testing by a pentanomial when $\deg{f}=\tilde{O}(\deg{g})$.
Theorem 1.18.
[…]
1.5 Proof Outline
We first note that Theorem 1.10 follows immediately from Theorem 1.8 and properties of the long division algorithm. The proof of Theorem 1.8 starts, similarly to [GR11, Theorem 2.9], by considering the discrete Fourier transform of $h$. By applying the transformation on the coefficient vector $h$ we obtain
$\left\|{h}\right\|_{2}^{2}=\frac{1}{p}\sum_{0\leq i<p}\left|h(\omega^{i})\right|^{2},$
where $\omega$ is a primitive $p$th root of unity. Thus, we conclude that $\left\|{h}\right\|_{2}\leq\left|h(\theta)\right|$ for some $p$th root of unity $\theta$. Since $\left|g(\theta)\right|\cdot\left|h(\theta)\right|=\left|f(\theta)\right|\leq\left\|{f}\right\|_{1}$, we get
$\left\|{h}\right\|_{2}\leq\left|h(\theta)\right|\leq\frac{\left\|{f}\right\|_{1}}{\left|g(\theta)\right|}.$
Hence, it is enough to prove an appropriate lower bound on the value of $g$ at roots of unity in order to obtain the result. To prove such a bound we consider the restriction of $g$ to the unit circle, $\tilde{g}(x)=g(e^{ix}):[0,2\pi)\to\mathbb{C}$. We prove by induction on the sparsity of $g$ that outside the neighborhood of a small set of bad points $B(g)\subset[0,2\pi)$, $g(e^{ix})$ attains large values. Indeed, when $\deg{g}=1$ this is relatively simple to show as $g$ must be of the form $x-\alpha$ (after rescaling) and since primitive roots of unity of relatively prime orders are somewhat far from each other, we can find such prime $p$ so that the values that $g$ gets on primitive $p$th roots of unity are not too small. For higher degrees, we note that if the derivative of $g$ is not small in a large enough region then $g$ is large on most of the region. Thus, by moving further away from the bad set for the derivative of $g$, $B(g^{\prime})$ we get that on this set $g$ always attains large values. We then prove by simple pigeonhole principle there exists a prime $p$ such that all primitive $p$th roots of unity are far away from the bad set and conclude the result.
The proof of Theorem 1.12 is based on two key ideas. First, we analyze the sparsity of a quotient of a division by a linear function. To do so, we use a special case of a proposition by Lenstra [Len99], which states that when $f$ can be written as $f=f_{0}+x^{d}f_{1}$ where $d-\deg{f_{0}}\geq\Omega(\log{\left\|{f}\right\|_{0}}+\log{\left\|{f}\right\|_{\infty}})$, then any zero of $f$ in $\mathbb{Q}$ that is not $1$, $-1$ nor 00 is a zero of both $f_{0}$ and $f_{1}$. We then observe that any binomial can be written as $g=x^{k}\ell(x^{m})$ for some linear function $\ell$ and $k,m\in\mathbb{N}$, and exploit this structure to conclude the bound on the sparsity of the quotient.
The proof of Theorem 1.13 is based on the same proposition by Lenstra for general roots in the algebraic closure of $\mathbb{Q}$. First, we prove a bound of the form $O\left(\left\|{f}\right\|_{0}\cdot\left(\frac{\deg{f}}{\deg{g}}\right)^{2}\right)$ on the sparsity of $h$ when $g$ is a trinomial, using an approach similar to [GGPdC21, Lemma 3.4]. We observe that trinomials have a large gap, either between the highest and second highest exponents or between the lowest and second lowest exponents. Thus, by analyzing $f/g$ over the ring of power series $\mathbb{F}[[x]]$, we conclude the result. Finally, we extend Lenstra’s approach to show that if $g$ is cyclotomic-free and $f$ can be split into two polynomials with a large degree gap, then $g$ divides $f$ if and only if $g$ divides both polynomials. This allows us to conclude that $h$ can be written as the sum of multiple quotients of polynomials whose degree ratio is smaller than $\deg{f}/\deg{g}$.
We leverage our bounds to conclude Theorem 1.14 and Theorem 1.15 by running the long division algorithm for polynomials with early termination in the case that the sparsity obtained exceeds the bound.
We prove Theorem 1.16 over finite fields by observing that for a binomial $g$, the remainder $f\bmod g$ must be sparse, and can be calculated explicitly. We prove Theorem 1.18 by first using a gap lemma for the case where $\deg{f}\leq\deg{g}(1+\frac{1}{\left\|{g}\right\|_{0}})$. In that case, $g$ must have a gap that is larger than $\deg{f}-\deg{g}$. This allows us to split $g$ into two parts, such that one of them has to be a binomial. We show that testing divisibility by this binomial is enough. Finally, we prove a reduction from testing the divisibility of a high-degree polynomial over a finite field to testing the divisibility of a polynomial of a higher sparsity but of a smaller degree, which yields a good trade-off when the field size is small. This allows us to finish the proof by showing that the case when $\deg{f}=\tilde{O}(\deg{g})$ can be reduced to the case $\deg{f}\leq\deg{g}(1+\frac{1}{\left\|{g}\right\|_{0}})$.
1.6 Open Problems
The main challenge is to extend our result over $\mathbb{Z}$ beyond trinomials and provide a polynomial time algorithm for every cyclotomic-free polynomial of bounded sparsity.
Another important question is to improve our result over $\mathbb{Z}$ to handle trinomials with cyclotomic factors, or to prove that this case is hard. We remark that our approach fails to do so only when $\deg{f}\geq\Omega(\deg{g}\log^{\omega(1)}\deg{g})$.
Another problem is extending our results over $\mathbb{F}_{p}$. For example, solving the case of $\deg{f}\geq\Omega(\deg^{1+\epsilon}{g})$ for pentanomials, extending the result to all bounded sparsities when $\deg{f}=\tilde{O}(\deg{g})$, or proving that either of those is hard are possible research directions.
Finally, we also note that the general question of the hardness of divisibility testing of lacunary polynomials (of variable sparsity) remains open, over both $\mathbb{Z}$ and $\mathbb{F}_{p}$.
1.7 Organization of the Paper
In Section 2 we set our notations and state useful facts and claims. In Section 3 we prove Theorem 1.8. In Section 4 we prove Theorems 1.12 and 1.13. In Section 5, we use the above bounds to prove Theorems 1.14 and 1.15. Finally, in Section 6 we deal with divisibility testing over finite fields and prove Theorems 1.16 and 1.18.
2 Preliminaries
In this section, we set our notation, and state some basic facts that we shall later use.
2.1 General Notation
Let $R$ be a unique factorization domain (UFD) and let $g,f\in R[x]$ be two polynomials. We say that $g$ divides $f$, and denote it with $g\mid f$, when there is a polynomial $h\in R[x]$ such that $g\cdot h=f$. We denote $g\nmid f$ when this is not the case.
Let $f=\sum_{i}f_{i}x^{i}$,. We shall use the following notation: $\left\|{f}\right\|_{0}$ denotes the sparsity of $f$ (the number of non-zero $f_{i}$’s); $\left\|{f}\right\|_{p}=\left(\sum_{i}{\left|f_{i}\right|^{p}}\right)^{1/p}$; if $R=\mathbb{Z}$ then $\left\|{f}\right\|_{\infty}=\max_{i}{\left|f_{i}\right|}$ is the height of $f$; $\text{ord}_{\alpha}{f}$ denotes the multiplicity of a root $\alpha$ in $f$, in particular, $\text{ord}_{0}{f}=\max{\left\{i:x^{i}\mid f\right\}}$; $\text{LC}(f)\in R$ denotes the leading coefficient in $f$, i.e., the coefficient of the highest degree term in $f$.
For any real or complex-valued function $f$, we define $V(f)$ to be the set of zeros of $f$, i.e., $V(f)=\left\{\alpha\mid f(\alpha)=0\right\}$. For a complex-valued function $f$ we denote by $\Re{f}$ and $\Im{f}$ the real and imaginary parts of $f$, respectively. We denote the derivative of $f$ by $f^{\prime}$.
A polynomial with $t$ non-zero terms is called a $t$-nomial; we call a polynomial with two terms a binomial, a polynomial with three terms a trinomial, and a polynomial with five terms a pentanomial.
For an integer $k\in\mathbb{N}$, we denote $[k]=\left\{1,\dots,k\right\}$. Throughout this paper, $\log$ refers to the natural logarithm. We use $\pi$ to denote the mathematical constant, while $\pi(\cdot)$ refers to the prime counting function. We use $i\;\equiv_{m}\;j$ to denote that $i=j\mod m$.
2.2 Useful Facts
2.2.1 Polynomials
An important class of polynomials, which contains the hardest instances for many problems concerning sparse polynomials, including all examples of NP-hard problems, is the class of cyclotomic polynomials. An $n$th root of unity, for a positive integer $n$, is any $\omega\in\mathbb{C}$ satisfying the equation $\omega^{n}=1$. It is said to be primitive if it is not a root of unity of any smaller order $m<n$. The minimal polynomials over $\mathbb{Q}$ of roots of unity are called cyclotomic polynomials. More explicitly, the $n$-th cyclotomic polynomial is defined as
$\Phi_{n}=\prod_{\begin{subarray}{c}1\leq k\leq n\\ \gcd{(k,n)}=1\end{subarray}}{(x-e^{2\pi i\frac{k}{n}})}=\prod_{\zeta}{(x-\zeta)},$
where $\zeta$ ranges over the primitive $n$th roots of unity. It is not hard to show that cyclotomic polynomials have integer coefficients. A cyclotomic-free polynomial is a polynomial that has no cyclotomic factors, and thus none of its zeros is a root of unity.
We next list some basic facts regarding polynomials.
Fact 2.1.
Let $R$ be a UFD, and let $f,g\in R[x]$. Then,
  1. 1.
    $g(x)\mid f(x)\iff g(x^{m})\mid f(x^{m})$ for any $0\neq m\in\mathbb{N}$.
  2. 2.
    $g\mid f\iff gh\mid fh$ for any $0\neq h\in R[x]$.
  3. 3.
    $g\mid f\iff g\mid fh$ for any $h\in R[x]$ such that $\gcd{\left\{g,h\right\}}=1$.
  4. 4.
    If $R$ is a field, then $g\mid f\iff g\mid af\iff ag\mid f$ for any $0\neq a\in R$.
Fact 2.2 (Polynomial Remainder Theorem).
Let $R$ be a commutative ring with a unit. Let $f\in R[x]$ and $a\in R$. Then $(x-a)\mid f\iff f(a)=0$.
Claim 2.3.
Let $R$ be a UFD. Let $g\in R[x]$ be a polynomial of sparsity $s$ and let $n\in\mathbb{N}$. Then $\left\|{g^{n}}\right\|_{0}\leq\binom{n+s-1}{s-1}$.
Proof.
Let $g=\sum_{i=1}^{s}g_{i}x^{d_{i}}$. Each monomial in $g^{n}$ is of the form $x^{\sum_{i=1}^{s}a_{i}d_{i}}$ where $\sum_{i=1}^{s}a_{i}=n$ and $a_{i}\geq 0$. Hence, each monomial in $g^{n}$ corresponds to at least one solution to the equation
$\sum_{i\in[s]}{a_{i}}=n$
over the non-negative integer variables $a_{1},\dots,a_{s}$, and there are exactly $\binom{n+s-1}{s-1}$ such solutions. ∎
Fact 2.4.
Let $g\in\mathbb{F}_{p}[x]$. Then $g^{p}=g(x^{p})$.
2.2.2 Complex Analysis
For general facts about complex analysis, see [Ahl21, Tit02]. For the definition of harmonic functions, see [Ahl21, Chapter 4.6] and [ABW13].
Claim 2.5.
Let $f:\mathbb{C}\to\mathbb{R}$ be a harmonic function. If $f(\omega)=0$ for every $\omega$ such that $\left|\omega\right|=1$, then $f=0$.
Proof.
Let,
$D=\left\{\omega\in\mathbb{C}:\left|\omega\right|\leq 1\right\}$
be the unit disc. From the Maximum Modulus Principle for harmonic functions (see [Ahl21, Chapter 4.6.2, Theorem 21]) and compactness of $D$ we get that $f(\omega)=0$ for all $\omega\in D$. It follows that $f$ must vanish everywhere [ABW13, Theorem 1.27]. ∎
Claim 2.6.
Let $f\in\mathbb{C}[x]$ be a polynomial. If $\Re{f}$ (or equivalently, $\Im{f}$, $\Re{f}+\Im{f}$ or $\Re{f}-\Im{f}$) is non-zero, then it has at most $2\deg{f}$ roots on the unit circle.
Proof.
Let $S\subseteq\mathbb{C}$ be the set of roots of $\Re{f}$ on the unit circle. We can write $\Re{f}(x+iy)=u(x,y)$ where $u:\mathbb{R}^{2}\to\mathbb{R}$ is a bi-variate polynomial of total degree $\deg{f}$. Hence,
$|S|=|V(u)\cap V(x^{2}+y^{2}-1)|.$
Since $f$ is analytic, $\Re{f}$ is harmonic. 2.5 implies that $\Re{f}$ does not vanish on the unit circle. Thus, from the irreducibility of $x^{2}+y^{2}-1$, we deduce that $V(u)$ and $V(x^{2}+y^{2}-1)$ have no component in common. From Bézout’s theorem (see [CLO13, Chapter 8.7,Theorem 10]) we conclude that
$|S|\leq\deg{u}\cdot\deg{(x^{2}+y^{2}-1)}=2\deg{f}.\qed$
Claim 2.7.
If $\left|x-y\right|\leq\pi$, then $\left|e^{ix}-e^{iy}\right|\geq\frac{2}{\pi}\left|x-y\right|$.
Proof.
Since
$\left|e^{ix}-e^{iy}\right|=\left|e^{i(x-y)}-1\right|,$
assume without loss of generality that $y=0$. If $\left|x\right|\leq\pi/2$, then
$\left|e^{ix}-1\right|=\sqrt{(1-\cos{x})^{2}+\sin^{2}{x}}\geq\left|\sin{x}\right|\geq\frac{2}{\pi}\left|x\right|,$
where the last inequality is due to the fact that $\sin$ is positive and concave on $[0,\pi]$, while negative and convex on $[-\pi,0]$. Similarly, if $\pi/2\leq\left|x\right|\leq\pi$, then
$\left|1-e^{ix}\right|=\sqrt{(1-\cos{x})^{2}+\sin^{2}{x}}\geq\left|1-\cos{x}\right|\geq\frac{2}{\pi}\left|x\right|,$
where the last inequality is due to the fact that $1-\cos$ is positive and concave on both $[\frac{\pi}{2},\frac{3\pi}{2}]$ and $[-\frac{3\pi}{2},-\frac{\pi}{2}]$. ∎
2.2.3 Number Theory
Fact 2.8 ([RS62, Corollary 1]).
Let $\pi(n)=\left|\left\{p\leq n:p\text{ is prime}\right\}\right|$ be the prime-counting function. Then $\pi(n)\geq\frac{n}{\log{n}}$ for $n\geq 17$.
Fact 2.9 ([RS62, Corollary 3]).
Let $n\in\mathbb{N}$ such that $n\geq 2$. Then, there exists at least one prime in $(n,2n]$.
Finally, we shall also use the following upper bound on binomial coefficients.
Claim 2.10.
Let $m,n\in\mathbb{N}$. Then, $\binom{n+m}{m}\leq(n+1)^{m}$.
Proof.
We prove this bound by induction on $n+m$. For the base case, assume $n+m=1$. Then, $\binom{n+m}{m}=1$ and the bound is trivial. Now assume that the bound is true for $n+m-1$. Then by Pascal’s triangle,
$\binom{n+m}{m}$ $=\binom{n+m-1}{m}+\binom{n+m-1}{m-1}$
$\leq n^{m}+(n+1)^{m-1}$
$\leq(n+1)^{m-1}n+(n+1)^{m-1}$
$=(n+1)^{m},$
which proves the inductive step. ∎
2.3 Preprocessing
In what follows, we assume that
  1. 1.
    $\text{ord}_{0}{g}=0$,
  2. 2.
    $\left\|{g}\right\|_{0}\geq 2$,
  3. 3.
    $\deg{g}\leq\deg{f}$, and
  4. 4.
    if $R$ is a field, then we also assume that $g$ and $f$ are monic.
To satisfy these assumptions, we need to introduce a preprocessing procedure to our algorithms. In particular, we can assume $\text{ord}_{0}{g}\leq\text{ord}_{0}{f}$ since $\text{ord}_{0}{g}>\text{ord}_{0}{f}$ implies that $g\nmid f$. Thus, we can divide both $f$ and $g$ by $x^{\text{ord}_{0}{g}}$ (see 2.1) and assume that $\text{ord}_{0}{g}=0$. If $g$ is a monomial, then $g\in R$, and so we can easily test $g\mid f$. If $\deg{g}>\deg{f}$ then $g\nmid f$. Finally, if $R$ is a field, the divisibility of $f$ by $g$ is invariant to multiplication by field elements (see 2.1). Thus we may assume that $g$ and $f$ are monic.
3 A Bound on the $\ell_{2}$-norm of the Quotient Polynomial
In this section, we prove Theorem 1.8. We do so by analyzing evaluations of polynomials at roots of unity of high orders.
The following lemma uses simple Fourier analysis to bound the norm of the quotient polynomial in a similar fashion to Theorem 2.9 of [GR11].
Lemma 3.1.
Let $f,g,h\in\mathbb{C}[x]$ such that $f=gh$. Then,
$\left\|{h}\right\|_{2}\leq\sqrt{2}\left\|{f}\right\|_{1}\cdot\max_{\begin{subarray}{c}1\neq\theta\in\mathbb{C}\\ \theta^{p}=1\end{subarray}}{\frac{1}{\left|g(\theta)\right|}}$
where $p$ is a prime such that $p>2\deg{f}$.
Proof.
Recall that the $p\times p$ Discrete Fourier Transform (DFT) matrix is
$\frac{1}{\sqrt{p}}\left(\omega^{jk}\right)_{0\leq j,k<p}$
where $\omega\in\mathbb{C}$ is a primitive $p$th root of unity. By applying the $p\times p$ DFT to the coefficient vector of $h$, we conclude that
$\left\|{h}\right\|_{2}^{2}=\sum_{0\leq i\leq\deg{h}}{\left|h_{i}\right|^{2}}=\frac{1}{p}\sum_{0\leq i<p}{\left|h(\omega^{i})\right|^{2}}$
where the second equality follows from the fact that the DFT matrix is unitary, and so it preserves $\ell_{2}$ norms. Observe that
$\left|h(\omega^{0})\right|^{2}=\left|h(1)\right|^{2}\leq\left\|{h}\right\|_{1}^{2}\leq\left\|{h}\right\|_{2}^{2}\deg{h}\leq\left\|{h}\right\|_{2}^{2}\deg{f},$
where the second inequality is due to Cauchy-Schwartz inequality. It follows that
$\frac{1}{p}\sum_{0<i<p}{\left|h(\omega^{i})\right|^{2}}=\left\|{h}\right\|_{2}^{2}-\frac{\left|h(\omega^{0})\right|^{2}}{p}\geq\left\|{h}\right\|_{2}^{2}-\frac{\deg{f}}{p}\left\|{h}\right\|_{2}^{2}>\frac{1}{2}\left\|{h}\right\|_{2}^{2},$
where the last inequality follows from the choice of $p$. In other words, the average value of $\left|h(\omega^{i})\right|^{2}$ for $0<i<p$ is larger than $\frac{1}{2}\left\|{h}\right\|_{2}^{2}$. Hence, there exists a $p$th root of unity $\theta\neq 1$ such that
$\left\|{h}\right\|_{2}<\sqrt{2}\left|h(\theta)\right|.$
Since $f=gh$ we conclude that $\left|f(\theta)\right|=\left|g(\theta)\right|\left|h(\theta)\right|$. If $g(\theta)=0$, then the claim trivially holds. As $\left|\theta\right|=1$, $\left|f(\theta)\right|\leq\left\|{f}\right\|_{1}$. Hence,
$\left\|{h}\right\|_{2}<\sqrt{2}\left|h(\theta)\right|=\sqrt{2}\frac{\left|f(\theta)\right|}{\left|g(\theta)\right|}\leq\sqrt{2}\frac{\left\|{f}\right\|_{1}}{\left|g(\theta)\right|}.$
By taking the maximum over all $p$-th roots of unity $\theta\neq 1$ we get
$\left\|{h}\right\|_{2}\leq\sqrt{2}\left\|{f}\right\|_{1}\cdot\max_{\begin{subarray}{c}1\neq\theta\in\mathbb{C}\\ \theta^{p}=1\end{subarray}}{\frac{1}{\left|g(\theta)\right|}}$
as claimed. ∎
To prove Theorem 1.8 we first study the case where $g$ is linear. We show that, in this case, there exists a point $\tilde{\alpha}$ on the unit circle such that $g$ attains large values on any point on the unit circle that is far enough from $\tilde{\alpha}$. To conclude, we prove that there exists a prime $p$ such that every $p$th root of unity is far from $\tilde{\alpha}$.
Lemma 3.2.
Let $\alpha_{1},\dots,\alpha_{k}\in\mathbb{R}$. Then, any set of $k+1$ prime numbers $P$ contains $p\in P$ such that $\left|a/p-\alpha_{i}\right|\geq\frac{1}{2\left(\max{P}\right)^{2}}$ for every integer $0<\left|a\right|<p$ and $i\in[k]$.
Proof.
Assume for the sake of contradiction that the statement is false. Thus, for every $p\in P$ there exists $i\in[k]$ such that $\left|a/p-\alpha_{i}\right|<\frac{1}{2\left(\max{P}\right)^{2}}$ for some $0<a<p$. By the Pigeonhole Principle, there exists $i\in[k]$ such that for two different primes $p,q\in P$, $0<a<p$ and $0<b<q$ we have
$\left|a/p-\alpha_{i}\right|<\frac{1}{2\left(\max{P}\right)^{2}}$    and $\left|b/q-\alpha_{i}\right|<\frac{1}{2\left(\max{P}\right)^{2}}.$
This, however, leads to a contradiction
$\frac{1}{\left(\max{P}\right)^{2}}>\left|\frac{a}{p}-\alpha_{i}\right|+\left|\alpha_{i}-\frac{b}{q}\right|\geq\left|\frac{a}{p}-\frac{b}{q}\right|=\left|\frac{ap-bq}{pq}\right|\geq 1/pq>\frac{1}{\left(\max{P}\right)^{2}}.$
Proposition 3.3.
Let $f\in\mathbb{C}[x]$ with a root $\alpha$. Then $\left\|{f/(x-\alpha)}\right\|_{2}\leq 64\sqrt{2}\left\|{f}\right\|_{1}\deg^{2}{f}<100\left\|{f}\right\|_{1}\deg^{2}{f}$.
Proof.
3.1 applied to $g=(x-\alpha)$ and $h=f/g$ implies that
$\left\|{f/(x-\alpha)}\right\|_{2}\leq\sqrt{2}\left\|{f}\right\|_{1}\cdot\max_{\begin{subarray}{c}1\neq\theta\in\mathbb{C}\\ \theta^{p}=1\end{subarray}}{\frac{1}{\left|\theta-\alpha\right|}},$ (3.4)
for any prime $p$ such that $p>2\deg{f}$. We prove that there exists a suitable $p$ such that $\left|\theta-\alpha\right|$ in Equation 3.4 must be at least $(64\deg^{2}{f})^{-1}$. First, observe that if $\alpha$ is $(64\deg^{2}{f})^{-1}$-far from any point on the unit circle, then we are done. So, assume that there exists a point $\tilde{\alpha}$ on the unit circle such that $\left|\alpha-\tilde{\alpha}\right|<(64\deg^{2}{f})^{-1}$, and define $c\in[-\pi,\pi)$ such that $e^{ic}=\tilde{\alpha}$.
2.9 implies that the interval $(2\deg{f},8\deg{f}]$ contains at least 2 primes. Thus, by 3.2, there exists a prime $p$ in this interval such that
$\left|\frac{2\pi a}{p}-c\right|=2\pi\left|\frac{a}{p}-\frac{c}{2\pi}\right|\geq\frac{\pi}{64\deg^{2}{f}},$
for any integer $0<\left|a\right|<p$. Observe that by the periodicity of $e^{ix}$, any $p$th root of unity $\theta\neq 1$ can be expressed as $\theta=e^{i\frac{2\pi a}{p}}$ such that $a\in\mathbb{Z}$, $0<\left|a\right|<p$ and $\left|\frac{2\pi a}{p}-c\right|\leq\pi$. Hence, by 2.7 it holds that
$\left|\theta-\tilde{\alpha}\right|\geq\frac{2}{\pi}\left|\frac{2\pi a}{p}-c\right|\geq\frac{1}{32\deg^{2}{f}},$
for any $p$th root of unity $\theta\neq 1$. The triangle inequality implies
$\left|\theta-\alpha\right|\geq\left|\theta-\tilde{\alpha}\right|-\left|\tilde{\alpha}-\alpha\right|\geq\frac{1}{64\deg^{2}{f}}.$
Finally, by using this estimate in Equation 3.4 we conclude that
$\left\|{f/(x-\alpha)}\right\|_{2}\leq\sqrt{2}\left\|{f}\right\|_{1}\cdot 64\deg^{2}{f}<100\left\|{f}\right\|_{1}\deg^{2}{f},$
which completes the proof. ∎
To generalize the above proposition, we need to establish a lower bound on $\left|g(\omega)\right|$ where $\omega\neq 1$ is a $p$th root of unity. We next show that $g$ attains large values everywhere, with the possible exception of a small neighborhood of a small number of points.
Claim 3.5.
Let $f:\mathbb{R}\to\mathbb{R}$ be a continuously differentiable function, $\delta>0$ and an interval $I=[a,b]$ in which each of $f$ and $f^{\prime}$ is either non-negative or non-positive. Then $|f(\beta)|\geq\delta\min_{\alpha\in I}|f^{\prime}(\alpha)|$ for every $\beta\in[a+\delta,b-\delta]$.
Proof.
If $b-a<2\delta$ there is nothing to prove. Assume without loss of generality that $f\geq 0$ in $I$, otherwise analyze $-f$. If $f^{\prime}\geq 0$ in $I$, then $f$ is non-decreasing in $I$ and so for every $\beta\in{[a+\delta,b-\delta]}$ it holds that
[…]
$\geq(\beta-a)\min_{\alpha\in I}|f^{\prime}(\alpha)|$
$\geq\delta\min_{\alpha\in I}|f^{\prime}(\alpha)|.$
Similarly, if $f^{\prime}\leq 0$ in $I$, then for every $\beta\in{[a+\delta,b-\delta]}$ it holds that
$f(\beta)\geq f(\beta)-f(b)$ $=\int_{\beta}^{b}{-f^{\prime}(y)dy}$
$\geq(b-\beta)\min_{\alpha\in I}|f^{\prime}(\alpha)|$
$\geq\delta\min_{\alpha\in I}|f^{\prime}(\alpha)|,$
as claimed. ∎
Claim 3.6.
Let $f:\mathbb{R}\to\mathbb{C}$ be a continuously differentiable function, $\delta>0$ and denote $f_{1}=\Re{f}$, $f_{2}=\Im{f}$. Let $I=[a,b]$ be an interval in which each function among $f_{1},f^{\prime}_{1},f_{2},f^{\prime}_{2},f_{1}^{\prime}-f_{2}^{\prime}$ and $f_{1}^{\prime}+f_{2}^{\prime}$ is either non-negative or non-positive. Then, $\left|f(\beta)\right|\geq\frac{\delta}{\sqrt{2}}\min_{\alpha\in I}|f^{\prime}(\alpha)|$ for every $\beta\in[a+\delta,b-\delta]$.
Proof.
Our assumption implies that $\left|f_{1}^{\prime}\right|-\left|f_{2}^{\prime}\right|$ is either non-negative or non-positive in $I$, as well. Without loss of generality assume $\left|f_{1}^{\prime}\right|\geq\left|f_{2}^{\prime}\right|$ in $I$. For every $\beta\in I$, $f^{\prime}(\beta)=f^{\prime}_{1}(\beta)+i\cdot f^{\prime}_{2}(\beta)$. Hence,
$2\left|f^{\prime}_{1}(\beta)\right|^{2}\geq\left|f^{\prime}_{1}(\beta)\right|^{2}+\left|f^{\prime}_{2}(\beta)\right|^{2}=\left|f^{\prime}(\beta)\right|^{2}\geq\min_{\alpha\in I}\left|f^{\prime}(\alpha)\right|^{2}.$
Thus, $\left|f_{1}^{\prime}(\beta)\right|\geq\frac{1}{\sqrt{2}}\min_{\alpha\in I}\left|f^{\prime}(\alpha)\right|$ for every $\beta\in I$. The result follows by 3.5 and the fact that $\left|f(\beta)\right|\geq\left|f_{1}(\beta)\right|$. ∎
In what comes next, we say that two points $\alpha$ and $\beta$ are $\delta$-far from each other if $\left|\alpha-\beta\right|\geq\delta$. Similarly, we say that a point is $\delta$-far from a set $S$ if it is $\delta$-far from any point in $S$.
Lemma 3.7.
Let $g\in\mathbb{C}[x]$ be a monic polynomial. Then, there exists a set $B(g)\subseteq[0,2\pi)$ of points, of size less than $12(\left\|{g}\right\|_{0}-1)\deg{g}$, such that for every $\delta>0$ and every $\alpha\in[0,2\pi)$ that is $(\left\|{g}\right\|_{0}-1)\delta$-far from $B(g)$ it holds that $\left|g(e^{i\alpha})\right|\geq\left(\frac{\delta}{\sqrt{2}}\right)^{\left\|{g}\right\|_{0}-1}$.
Proof.
Observe that $\left|(g/x^{\text{ord}_{0}{g}})(\omega)\right|=\left|g(\omega)\right|$ for every $\omega$ on the unit circle and so we can assume without loss of generality that $\text{ord}_{0}{g}=0$. We prove this lemma by induction on $s=\left\|{g}\right\|_{0}$. The claim holds for $s=1$ since it implies that $g=1$, and so clearly $|g(e^{i\alpha})|=1$ for any $\alpha$. For the inductive step, suppose that the claim holds for $s$, and let $g\in\mathbb{C}[x]$ be of sparsity $s+1$. Let $f:[0,2\pi)\to\mathbb{C}$ defined as $f(x)=g(e^{ix})$. Let,
$H$ =$=\left\{\Re f,\Re f^{\prime},\Im f,\Im f^{\prime},\Re f^{\prime}-\Im f^{\prime},\Re f^{\prime}+\Im f^{\prime}\right\}\text{ and }$ (3.8)
$B(g)$ $=B(g^{\prime}/(\deg{g}\cdot x^{\text{ord}_{0}{g^{\prime}}}))\cup\left\{\alpha\in[0,2\pi):h(\alpha)=0\text{ for a non-zero }h\in H\right\}.$
By 2.6, each non-zero function in $H$ has at most $2\deg{g}$ zeros in $[0,2\pi)$. Since $g^{\prime}/(\deg{g}\cdot x^{\text{ord}_{0}{g^{\prime}}})$ is clearly of sparsity $s$, by the inductive assumption we conclude that
$\left|B(g)\right|<12(s-1)\deg{g^{\prime}}+12\deg{g}<12s\deg{g}.$
Let $I=[a,b]\subseteq[0,2\pi)$ be an interval that is $(s-1)\delta$-far from $B(g)$. By definition of $B(g)$, each of the functions in the set $H$ defined in Equation (3.8) does not change its sign in $I$ (as every sign change would have to go through a zero of the function). As $f^{\prime}(x)=ie^{ix}\cdot{g^{\prime}}(e^{ix})$, 3.6 implies that for every $\beta\in[a+\delta,b-\delta]$ it holds that
$\left|f(\beta)\right|$ $\geq\frac{\delta}{\sqrt{2}}\min_{\alpha\in I}\left|f^{\prime}(\alpha)\right|=\frac{\delta}{\sqrt{2}}\min_{\alpha\in I}\left|g^{\prime}(e^{i\alpha})\right|=\frac{\delta}{\sqrt{2}}\min_{\alpha\in I}\left|(g^{\prime}/x^{\text{ord}_{0}{g^{\prime}}})(e^{i\alpha})\right|$
$\geq\frac{\delta}{\sqrt{2}}\min_{\alpha\in I}\left|(g^{\prime}/(\deg{g}\cdot x^{\text{ord}_{0}{g^{\prime}}}))(e^{i\alpha})\right|.$
As $B(g^{\prime}/(\deg{g}\cdot x^{\text{ord}_{0}{g^{\prime}}}))\subseteq B(g)$ and each $\alpha\in I$ is at least $(s-1)\delta$-far from $B(g)$, we get from the induction hypothesis applied to $g^{\prime}/(\deg{g}\cdot x^{\text{ord}_{0}{g^{\prime}}})$ that $\left|f(\beta)\right|\geq\left(\frac{\delta}{\sqrt{2}}\right)^{s}$. In other words, every $\beta$ that is $s\delta$ far from the set $B(g)$ satisfies $\left|f(\beta)\right|\geq\left(\frac{\delta}{\sqrt{2}}\right)^{s}$, which proves the step of the induction. ∎
Lemma 3.9.
Let $g\in\mathbb{C}[x]$ be a monic polynomial. Let $p_{min}\in\mathbb{N}$ and $L(n)=n\log{n}$. Set $p_{max}=\lceil 2L(p_{min}+12\left\|{g}\right\|_{0}\deg{g})\rceil$. Then, there exists a prime $p\in(p_{min},p_{max}]$ such that
$\left|g(\omega)\right|\geq\left(\frac{\pi}{\sqrt{2}\left\|{g}\right\|_{0}p_{max}^{2}}\right)^{\left\|{g}\right\|_{0}-1}\geq\left(\frac{2}{\left\|{g}\right\|_{0}p_{max}^{2}}\right)^{\left\|{g}\right\|_{0}-1},$
for any $p$th root of unity $\omega\neq 1$.
Proof.
If $\left\|{g}\right\|_{0}=1$, then the statement is trivial since $\left|g(\omega)\right|=1$ for every root of unity $\omega$. Hence, we can assume that $\left\|{g}\right\|_{0}\geq 2$ and $\deg{g}\geq 1$.
Let $t=p_{min}+12\left\|{g}\right\|_{0}\deg{g}$, and let $\pi(n)$ be the prime counting function. Our assumption implies that $t\geq 24$. By 2.8 we conclude that
$\pi(p_{max})$ $=\pi(\lceil 2t\log{t}\rceil)$
$\geq\frac{2t\log{t}}{\log{2}+\log{t}+\log{\log{t}}}$
$>\frac{2t\log{t}}{2\log{t}}=t.$
Let $B(g)$ be as in 3.7 and
$P=\left\{p_{min}<p\leq p_{max}:p\text{ is prime}\right\}.$
We bound the size of $P$ by counting the primes in the interval $(p_{min},p_{max}]$:
$\left|P\right|\geq\pi(p_{max})-p_{min}\geq t-p_{min}=12\left\|{g}\right\|_{0}\deg{g}>\left|B(g)\right|.$
By 3.2, there exists $p\in P$ such that for any integer $0<a<p$ and $\alpha\in B(g)$
$\left|a/p-\alpha/2\pi\right|\geq\frac{1}{2p_{max}^{2}}.$
Thus,
$\left|2\pi\frac{a}{p}-\alpha\right|\geq\frac{\pi}{p_{max}^{2}}.$
Finally, by applying 3.7 on $\delta=\frac{\pi}{\|g\|_{0}p_{max}^{2}}$, we conclude that for any $p$th root of unity $\omega\neq 1$,
$\left|g(\omega)\right|\geq\left(\frac{\pi}{\sqrt{2}\left\|{g}\right\|_{0}p_{max}^{2}}\right)^{\left\|{g}\right\|_{0}-1}.\qed$
Theorem 3.10.
Let $f,g,h\in\mathbb{C}[x]$ such that $f=gh$. Denote $T(n)=n\log{n}$. Then,
$\left\|{h}\right\|_{2}\leq\sqrt{2}\frac{\left\|{f}\right\|_{1}}{LC(g)}\left(3\left\|{g}\right\|_{0}\cdot L\left(12\left\|{g}\right\|_{0}\deg{g}+2\deg{f}\right)^{2}\right)^{\left\|{g}\right\|_{0}-1}.$
Proof.
First, we can assume that $g$ is monic: if $g$ is not monic, then we scale both $f$ and $g$ by $\frac{1}{LC(g)}$. Let $p_{min}=2\deg{f}$ and $p_{max}=\lceil 2L(12\|g\|_{0}\deg{g}+2\deg{f})\rceil$. By plugging those values into 3.9 we conclude that there exists a prime $p\in(p_{min},p_{max}]$ such that
$\left|g(\omega)\right|\geq\left(\frac{2}{\left\|{g}\right\|_{0}\cdot(\lceil 2L(12\left\|{g}\right\|_{0}\deg{g}+2\deg{f})\rceil)^{2}}\right)^{\left\|{g}\right\|_{0}-1}$
for any primitive $p$th root of unity $\omega\neq 1$. The theorem then follows from 3.1. ∎
We note that when $g$ does not divide $f$, no such bound holds for the remainder.
Remark 3.11.
Let $f,g\in\mathbb{C}[x]$ and let $q,r$ be the quotient and the remainder of $f$ divided by $g$, respectively. For any $\left\|{f}\right\|_{\infty}$, $\left\|{g}\right\|_{\infty}$, $\deg{f}$, and $\deg{g}$, there exists $f$ and $g$ of sparsity $2$ such that $\left\|{r}\right\|_{\infty}=\left\|{f}\right\|_{\infty}\left\|{g}\right\|_{\infty}^{\deg{f}-\deg{g}+1}$.
Proof.
Let $d_{1},d_{2}\in\mathbb{N}$ and $a\in\mathbb{C}$ with $\left|a\right|\geq 1$. Then,
$x^{d_{1}}\bmod x^{d_{2}+1}-ax^{d_{2}}=a^{d_{1}-d_{2}}x^{d_{2}}\qed$
The proof of Theorem 1.10 follows immediately from the bound proved in Theorem 3.10 and the analysis of the long division algorithm given in Section 5.
4 A Bound on the Sparsity of the Quotient Polynomial
Let $f,g,h\in\mathbb{Z}[x]$ such that $f=gh$. Throughout this section we assume, without loss of generality, that $\text{ord}_{0}{g}=0$. Indeed, when this does not hold, we divide $f$ and $g$ by $x^{\text{ord}_{0}{g}}$. This does not change the coefficient sizes and sparsities of $f$ and $g$, and it only reduces their degrees.
4.1 Division by Binomial
In this section, we prove Theorem 1.12. We bound the sparsity of $h$ in terms of $f$ and $g$ when $g$ is a cyclotomic-free binomial. We first deal with the case when $g$ is a linear function. The following lemma is a special case of a result of Lenstra [Len99] that plays an important role in our proofs.
Lemma 4.1 (Proposition 2.3 of [Len99] for the linear case).
Let $f\in\mathbb{Z}[x]$ such that $f=f_{0}+x^{d}f_{1}$. Suppose that
$d-\deg{f_{0}}\geq\frac{1}{\log{2}}\left(\log{(\left\|{f}\right\|_{0}-1)}+\log{\left\|{f}\right\|_{\infty}}\right).$
Then, every zero of $f$ in $\mathbb{Q}$ that is not $1$, $-1$ nor 00 is a common zero of both $f_{0}$ and $f_{1}$.
The next argument is similar to the one in [GRT12, Lemma 4.4].
Lemma 4.2.
Let $f,\ell,h\in\mathbb{Z}[x]$ such that $f=\ell h$ and $\ell$ is linear. If $1$ and $-1$ are not roots of $\ell$, then the sparsity of $h$ is at most
$\frac{1}{\log{2}}\left\|{f}\right\|_{0}\left(\log{(\left\|{f}\right\|_{0}-1)}+\log{\left\|{f}\right\|_{\infty}}\right).$
Proof.
Assume without loss of generality that $\text{ord}_{0}{\ell}=0$. Let $\alpha$ be the only root of $\ell$ and observe that it is clearly in $\mathbb{Q}$. Find a decomposition of $f$ to polynomials with large gaps between their monomials,
$f=f_{0}+\sum_{i\in[k]}{x^{d_{i}}f_{i}},$
such that
$d_{i}-d_{i-1}-\deg{f_{i-1}}\geq\frac{1}{\log{2}}\left(\log{(\left\|{f}\right\|_{0}-1)}+\log{\left\|{f}\right\|_{\infty}}\right),$
for all $i\in[k]$, and each $f_{i}$ cannot be further decomposed in the same way. Such a decomposition exists by a simple greedy algorithm and is unique. As $\left\|{f_{i-1}}\right\|_{0}\leq\left\|{f}\right\|_{0}$ and $\left\|{f_{i-1}}\right\|_{\infty}\leq\left\|{f}\right\|_{\infty}$ we conclude that
$d_{i}-d_{i-1}-\deg{f_{i-1}}\geq\frac{1}{\log{2}}\left(\log{(\left\|{f_{i-1}}\right\|_{0}-1)}+\log{\left\|{f_{i-1}}\right\|_{\infty}}\right).$
By applying 4.1 recursively we conclude that $f_{i}(\alpha)=0$ for all $0\leq i\leq k$. Observe that as $f_{i}$ cannot be further decomposed, it holds that
$\frac{\deg{f_{i}}}{\left\|{f_{i}}\right\|_{0}}<\frac{1}{\log{2}}\left(\log{(\left\|{f}\right\|_{0}-1)}+\log{\left\|{f}\right\|_{\infty}}\right),$
Hence,
$\left\|{h}\right\|_{0}$ $\leq\sum_{i=0}^{k}{\left\|{\frac{f_{i}}{\ell}}\right\|_{0}}$
$\leq\sum_{i=0}^{k}{\deg{\frac{f_{i}}{\ell}}}$
$\leq\sum_{i=0}^{k}{\deg{{f_{i}}}}$
$\leq\sum_{i=0}^{k}{\frac{1}{\log{2}}\left\|{f_{i}}\right\|_{0}\left(\log{(\left\|{f}\right\|_{0}-1)}+\log{\left\|{f}\right\|_{\infty}}\right)}$
$\leq\frac{1}{\log{2}}\left\|{f}\right\|_{0}\left(\log{(\left\|{f}\right\|_{0}-1)}+\log{\left\|{f}\right\|_{\infty}}\right),$
which completes the proof. ∎
Remark 4.3.
The requirement that $g$ is cyclotomic-free is essential, as demonstrated by the following examples,
$\frac{x^{m}-1}{x-1}$ $=\sum_{0\leq i<m}{x^{i}}$
$\frac{x^{m}+1}{x+1}$ $=\sum_{0\leq i<m}{(-x)^{i}},$
where $m\in\mathbb{N}$ is odd.
Remark 4.4 (Tightness of 4.2).
For any integer value of $\log_{2}{\left\|{f}\right\|_{\infty}}$ and $\left\|{f}\right\|_{0}>1$, there exists $f,g,h\in\mathbb{Z}[x]$ as in Theorem 4.6 with $h$ having sparsity $(\left\|{f}\right\|_{0}-1)\log_{2}{\left\|{f}\right\|_{\infty}}$. For example,
$\frac{x^{k+1}-2^{k+1}}{x-2}=\sum_{i=0}^{k}{2^{k-i}x^{j}}$
for any $k\in\mathbb{N}$. Letting $f=\sum_{j=0}^{m}{x^{kj}(x^{k}-2^{k})}$ and $g=x-2$ yields such an example.
To generalize 4.2 to binomials of any degree, we observe that any binomial $g$ can be written as $g=x^{k}\ell(x^{m})$ for some linear function $\ell$ and $m,k\in\mathbb{N}$. The following lemma deals with polynomials of this form, and is used in other contexts later in the paper.
Lemma 4.5.
Let $R$ be a UFD, and let $m\in\mathbb{N}$. Let $f,\ell\in R[x]$. Denote $f=\sum_{i}{f_{i}x^{i}}$ and let
$u_{j}:=\sum_{i\;\equiv_{m}\;j}{f_{i}x^{(i-j)/m}}.$
Then, $\ell(x^{m})\mid f$ if and only if $\ell\mid u_{j}$ for all $0\leq j<m$, and in this case we have that
$h=\sum_{0\leq j<m}{x^{j}v_{j}(x^{m})},$
where $h=\frac{f}{\ell(x^{m})}$ and $v_{j}=\frac{u_{j}}{\ell}$.
Proof.
First, observe that $f=\sum_{0\leq j<m}{x^{j}u_{j}(x^{m})}$. For the first direction, assume that $u_{j}=\ell v_{i}$ for all $0\leq i<m$. Then it holds that $u_{j}(x^{m})=\ell(x^{m})v_{i}(x^{m})$. Thus $x^{j}u_{j}(x^{m})=\ell(x^{m})\cdot x^{j}v_{i}(x^{m})$, and we conclude that
$f=\sum_{0\leq j<m}{x^{j}u_{j}(x^{m})}=\ell(x^{m})\sum_{0\leq j<m}{x^{j}v_{i}(x^{m})}=\ell(x^{m})h$
as needed. To prove the converse, assume that $f=\ell(x^{m})h$ and define $h=\sum_{i}{h_{i}x^{i}}$. Then,
$f$ $=\ell(x^{m})\left(\sum_{i}{h_{i}x^{i}}\right)$
$=\ell(x^{m})\left(\sum_{0\leq j<m}{\left(\sum_{i\;\equiv_{m}\;j}{h_{i}x^{i}}\right)}\right)$
$=\sum_{0\leq j<m}{\left(\ell(x^{m})\sum_{i\;\equiv_{m}\;j}{h_{i}x^{i}}\right)}.$
By comparing exponents we conclude that for all $0\leq j<m$,
$\sum_{i\;\equiv_{m}\;j}{f_{i}x^{i}}=\ell(x^{m})\sum_{i\;\equiv_{m}\;j}{h_{i}x^{i}}.$
Hence,
$\sum_{i\;\equiv_{m}\;j}{f_{i}x^{i-j}}=\ell(x^{m})\sum_{i\;\equiv_{m}\;j}{h_{i}x^{i-j}},$
which implies that
$u_{j}=\sum_{i\;\equiv_{m}\;j}{f_{i}x^{(i-j)/m}}=\ell\sum_{i\;\equiv_{m}\;j}{h_{i}x^{(i-j)/m}}=\ell v_{j},$
as claimed. ∎
We are ready to conclude the sparsity bound for dividing by a binomial, by reducing the problem to the linear case.
Theorem 4.6.
Let $f,g,h\in\mathbb{Z}[x]$ such that $f=gh$ and $g$ is a binomial. If $g$ is cyclotomic-free, then $h$ has sparsity of at most
$\frac{1}{\log{2}}\left\|{f}\right\|_{0}\left(\log{(\left\|{f}\right\|_{0}-1)}+\log{\left\|{f}\right\|_{\infty}}\right).$
Proof.
Assume without loss of generality that $\text{ord}_{0}{g}=0$. Let $m\in\mathbb{N}$ be such that $g=\ell(x^{m})$ for a linear $\ell$. For $u_{j}$ as defined in 4.5 we get that
$h=\sum_{0\leq j<m}{x^{j}\left(\frac{u_{j}(x^{m})}{\ell(x^{m})}\right)}.$
From 4.2 we get
$\left\|{h}\right\|_{0}$ $\leq\sum_{0\leq j<m}{\left\|{\frac{u_{j}}{\ell}}\right\|_{0}}$
$\leq\frac{1}{\log{2}}\sum_{0\leq j<m}{\left\|{u_{j}}\right\|_{0}\left(\log{(\left\|{u_{j}}\right\|_{0}-1)}+\log{\left\|{f}\right\|_{\infty}}\right)}$
$\leq\frac{1}{\log{2}}\left\|{f}\right\|_{0}\left(\log{(\left\|{f}\right\|_{0}-1)}+\log{\left\|{f}\right\|_{\infty}}\right).\qed$
4.2 Division by Trinomial
In this section, we prove Theorem 1.13. We bound the sparsity of $h$ in terms of $f$ and $g$ when $g$ is a cyclotomic-free trinomial. To achieve this, we first bound $h$ when $\deg{f}=\tilde{O}(\deg{g})$. We then incorporate a proposition by Lenstra to achieve a bound for the general case.
4.2.1 The Case of $\deg{f}=\tilde{O}(\deg{g})$
We show that in this case the sparsity of $h=f/g$ is upper bounded by $2\left\|{f}\right\|_{0}\cdot\left(\frac{\deg{f}}{\deg{g}}\right)^{2}$ (4.8). The proof is similar to the proof in [GGPdC21, Section 3]. We give a different analysis that better suits our purpose.
Claim 4.7.
Let $R$ be an integral domain. Let $f,g,h\in R[x]$ such that $f=gh$ and $\text{ord}_{0}{g}=0$. If $g=c-x^{d}\hat{g}$ for $c\in R$, then $h$ has sparsity at most
$\left\|{f}\right\|_{0}\cdot\sum_{k=0}^{\left\lfloor\frac{\deg{f}-\deg{g}}{d}\right\rfloor}{(k+1)^{\left\|{g}\right\|_{0}-2}}.$
Proof.
Let $\mathbb{F}$ be the field of fractions of $R$, and define $\tilde{g}\in\mathbb{F}[x]$ such that $\tilde{g}=c^{-1}\hat{g}$. We analyze $h$ over the ring of power series $\mathbb{F}[[x]]$. Since $\frac{1}{1-\tilde{g}}=\sum_{k=0}^{\infty}{\tilde{g}^{k}}$ we conclude that
$\frac{1}{g}=\frac{1}{c(1-x^{d}\tilde{g})}=c^{-1}\cdot\sum_{k=0}^{\infty}{\tilde{g}^{k}x^{dk}}.$
Since $f=gh$, it holds that $\deg{h}=\deg{f}-\deg{g}$. Hence,
$h$ $=\frac{f}{g}$
$=c^{-1}f\cdot\sum_{k=0}^{\infty}{\tilde{g}^{k}x^{dk}}\mod x^{\deg{f}-\deg{g}+1}$
$=c^{-1}f\cdot\sum_{k=0}^{\left\lfloor\frac{\deg{f}-\deg{g}}{d}\right\rfloor}{\tilde{g}^{k}x^{dk}}\mod x^{\deg{f}-\deg{g}+1}.$
Finally, by comparing monomials and applying 2.3 and 2.10 we conclude that
$\left\|{h}\right\|_{0}$ $\leq\left\|{f}\right\|_{0}\cdot\sum_{k=0}^{\left\lfloor\frac{\deg{f}-\deg{g}}{d}\right\rfloor}{\left\|{\tilde{g}^{k}}\right\|_{0}}$
$\leq\left\|{f}\right\|_{0}\cdot\sum_{k=0}^{\left\lfloor\frac{\deg{f}-\deg{g}}{d}\right\rfloor}{(k+1)^{\left\|{\tilde{g}}\right\|_{0}-1}},$
which proves the claim. ∎
The next lemma shows that when $\deg{f}=\tilde{O}(\deg{g})$ and $f=gh$, the sparsity of $h$ is polynomial in the size of the representation.
Lemma 4.8.
Let $R$ be an integral domain. Let $f,g,h\in R[x]$ such that $f=gh$ and $g$ is a trinomial. Then, $h$ has sparsity at most
$2\left\|{f}\right\|_{0}\cdot\left(\frac{\deg{f}}{\deg{g}}\right)^{2}.$
Proof.
First, assume without loss of generality that $\text{ord}_{0}{g}=0$. Let $d$ as in 4.7. Assume without loss of generality that $d\geq\frac{1}{2}\deg{g}$. Indeed, if $d<\frac{1}{2}\deg{g}$ then consider the polynomials $\phi(f),\phi(g),\phi(h)$, where $\phi:R[x]\to R[x]$ is defined as $\phi(f)=x^{\deg{f}}f(1/x)$. Observe that $\phi(f)=\phi(g)\phi(h)$ and that $\phi$ preserves sparsities (see remark 3.2 of [GGPdC21]). By 4.7 it holds that
$\left\|{h}\right\|_{0}$ $\leq\left\|{f}\right\|_{0}\cdot\sum_{k=0}^{\left\lfloor\frac{\deg{f}-\deg{g}}{d}\right\rfloor}{(k+1)^{\left\|{g}\right\|_{0}-2}}$
$\leq\left\|{f}\right\|_{0}\cdot\sum_{k=0}^{\left\lfloor 2(\frac{\deg{f}}{\deg{g}}-1)\right\rfloor}{(k+1)}$
$\leq\left\|{f}\right\|_{0}\cdot\frac{\deg{f}}{\deg{g}}\left(2\frac{\deg{f}}{\deg{g}}-1\right)$
$\leq 2\left\|{f}\right\|_{0}\cdot\left(\frac{\deg{f}}{\deg{g}}\right)^{2},$
as claimed. ∎
Remark 4.9 (Tightness of 4.8).
For any value of $\deg{g}$, $\deg{f}$, and $\left\|{f}\right\|_{0}>1$, there exist $f,g,h\in\mathbb{Z}[x]$ as in 4.8 with $h$ having sparsity $O\left(\left\|{f}\right\|_{0}\frac{\deg{f}}{\deg{g}}\right)$. For example,
$\frac{x^{6k}-1}{x^{2}+x+1}=\sum_{i=0}^{2k-1}{x^{3i}(x-1)}$
for any $k\in\mathbb{N}$. Taking $f=\sum_{j=0}^{m}{x^{6knj}(x^{6kn}-1)}$ and $g=x^{2n}+x^{n}+1$ for some $n,m\in\mathbb{N}$ proves the remark.
Remark 4.10 (Division by Tetranomials).
For any values of $\deg{f}$ and $\deg{g}$, there exists $f,g,h\in\mathbb{Z}[x]$ such that $f=gh$ and $f$ and $g$ are tetranomials, but $h$ has sparsity $\deg{f}-\deg{g}+1$. For example,
$\frac{x^{k+n}-x^{n}+x^{k}-1}{x^{n+1}-x^{n}+x-1}=\sum_{i=0}^{k-1}{x^{i}}$
for any $n,k\in\mathbb{N}$.
Remark 4.11 (Comparison to [GGPdC21]).
When restricted to trinomials, Corollaries 3.4 and 3.5 of [GGPdC21] imply a bound that is exponential in $\frac{\deg{f}}{\deg{g}}$. Thus, our analysis is tighter in this case.
4.2.2 The Case of Cyclotomic-Free Trinomials
To obtain the sparsity bounds we use a lemma that is based on a work of Lenstra [Len99]. In what follows, $\overline{\mathbb{Q}}\subset\mathbb{C}$ denotes the algebraic closure of $\mathbb{Q}$. We say that an algebraic number $\alpha\in\overline{\mathbb{Q}}$ is of degree $n$ over $\mathbb{Q}$ if the monic polynomial of least degree $\ell\in\mathbb{Q}[x]$ such that $\ell(\alpha)=0$ is of degree $n$.
Lemma 4.12 (Proposition 2.3 of [Len99]).
Let $f\in\mathbb{Z}[x]$ such that $f=f_{0}+x^{d}f_{1}$. Suppose that $t\geq 2$ is a positive integer with
$d-\deg{f_{0}}\geq\frac{1}{2}t\log^{3}{(3t)}\left(\log{(\left\|{f}\right\|_{0}-1)}+\log{\left\|{f}\right\|_{\infty}}\right).$
Then, every zero of $f$ that has a degree at most $t$ over $\mathbb{Q}$ and that is not a root of unity nor zero is a common zero of both $f_{0}$ and $f_{1}$.
We aim to show that, under similar assumptions, if $g$ is a factor of $f_{0}+x^{d}f_{1}$, then $g$ is a factor of both $f_{0}$ and $f_{1}$. For that we extend Lenstra’s proposition to take into account the multiplicities of zeros of $f$ as well.
Lemma 4.13.
Let $f\in\mathbb{Z}[x]$ such that $f=f_{0}+x^{d}f_{1}$. Suppose that $t\geq 2$ and $m$ are positive integers with
$d-\deg{f_{0}}\geq\frac{1}{2}t\log^{3}{(3t)}\left(\log{\left\|{f}\right\|_{\infty}}+m(5+3\log{\deg{f}})\right).$
Then, every root of $f$, which is not a root of unity nor zero, that has multiplicity $m$ and degree at most $t$ over $\mathbb{Q}$ is a common zero of both $f_{0}$ and $f_{1}$ of the same multiplicity.
Proof.
Let $\alpha$ be a root of $f$ of degree $t$ and multiplicity $m$. We prove the lemma by induction. If $m=1$, apply 4.12 to $f$ and conclude the result. Now assume the statement is true for $m-1$, and let $g=f/(x-\alpha)$. Clearly,
  1. 1.
    The multiplicity of $\alpha$ in $g$ is exactly $m-1$.
  2. 2.
    $\deg{g}\leq\deg{f}$
  3. 3.
    Denote $g_{0}=f_{0}/(x-\alpha)$ and $g_{1}=f_{1}/(x-\alpha)$, and observe that $d-\deg{g_{0}}\geq d-\deg{f_{0}}$.
  4. 4.
    By 3.3, it holds that
    $\left\|{g}\right\|_{\infty}\leq\left\|{g}\right\|_{2}\leq 100\deg^{2}{f}\left\|{f}\right\|_{1}\leq 100\deg^{3}{f}\left\|{f}\right\|_{\infty}$
Hence, $g$ has a gap of size,
$d-\deg{g_{0}}$ $\geq\frac{1}{2}t\log^{3}{(3t)}\left(\log{\left\|{f}\right\|_{\infty}}+m(5+3\log{\deg{f}}\right)$
$\geq\frac{1}{2}t\log^{3}{(3t)}\left(\log{\left\|{g}\right\|_{\infty}}+(m-1)(5+3\log{\deg{g}}\right).$
By the inductive assumption, this means that $\alpha$ is a root of both $g_{0}$ and $g_{1}$ of multiplicity $m-1$, and so it is a root of both $f_{0}$ and $f_{1}$ of multiplicity $m$. ∎
Claim 4.14.
Let $f\in\mathbb{C}[x]$. Then, any nonzero root of $f$ has multiplicity $\leq\left\|{f}\right\|_{0}-1$.
Proof.
See for example Lemma 1 of [GKO92]. ∎
Corollary 4.15.
Let $f,g\in\mathbb{Z}[x]$ such that, $\text{ord}_{0}{g}=0$ and $g$ is cyclotomic-free. Denote $f=f_{0}+x^{d}f_{1}$ such that $d-\deg{f_{0}}\geq b(f,g)$ where
$b(f,g)=\frac{1}{2}\deg{g}\log^{3}{(3\deg{g})}\left(\log{\left\|{f}\right\|_{\infty}}+(\left\|{g}\right\|_{0}-1)(5+3\log{\deg{f}})\right).$
Then, $g\mid f$ if and only if $g\mid f_{0}$ and $g\mid f_{1}$.
Proof.
Let $\alpha$ be a root of $g$. Then, its degree is bounded by $\deg{g}$, and by 4.14, its multiplicity is bounded by $\left\|{g}\right\|_{0}$ 1. By 4.13, we conclude that $\alpha$ is a root of $f$ if and only if it is a root of both $f_{0}$ and $f_{1}$ of the same multiplicity. The claim follows as this is true for every root of $g$. ∎
Remark 4.16 (Testing Divisibility by a Low-degree Polynomial over $\mathbb{Z}$).
Unlike the case of finite fields, where the remainder can be directly computed, testing divisibility by a low degree polynomial over $\mathbb{Z}[x]$ is not entirely trivial as the remainder can have exponentially (in the degree) large coefficients (e.g. this is the case when $f=x^{n}$ and $g=x-2$) and thus cannot be explicitly computed in polynomial time.
However, 4.15 gives rise to an efficient divisibility testing algorithm: we can recursively decompose the polynomial $f$ into $f_{0},\dots,f_{k}$, and test $g\mid f_{i}$ by calculating the remainder directly for all $0\leq i\leq k$. See the proof of Theorem 4.17 for how to find such a decomposition.
Alternatively, such a test can be obtained by factoring $g$, and then applying Lenstra’s algorithm [Len99] for finding low-degree factors of $f,f^{\prime},\dots,f^{(m)}$ for $m=\left\|{g}\right\|_{0}-1$.
Finally, recall that we started the discussion with $f,g,h\in\mathbb{Z}[x]$ such that $f=gh$ and $g$ is a trinomial. 4.15 allows us to bound the sparsity of $h$ in terms of $f,g$ when $g$ is also cyclotomic-free.
Theorem 4.17.
Let $f,g,h\in\mathbb{Z}[x]$ such that $f=gh$ and $g$ is a cyclotomic-free trinomial. Then, the sparsity of $h$ is at most
$\frac{1}{2}\left\|{f}\right\|_{0}^{3}\log^{6}{(3\deg{g})}\left(\log{\left\|{f}\right\|_{\infty}}+6\log{\deg{f}}+10\right)^{2}.$
Proof.
First, assume without loss of generality that $\text{ord}_{0}{g}=0$. Let $b(f,g)$ as in 4.15. Define,
$f=f_{0}+\sum_{i\in[k]}{x^{d_{i}}f_{i}},$
such that
$d_{i}-d_{i-1}-\deg{f_{i-1}}\geq b(f,g),$
for all $i\in[k]$, and each $f_{i}$ cannot be further decomposed in the same way. By applying 4.15 recursively we conclude that $g\mid f_{i}$ for all $0\leq i\leq k$ (this can be done since the sparsity, degree, and coefficients of $f_{i}$ are bounded by those of $f$). Observe that for all $0\leq i\leq k$ it holds that
$\frac{\deg{f_{i}}}{\left\|{f_{i}}\right\|_{0}}<b(f,g),$
since it cannot be further decomposed. Hence, by 4.8 we conclude that
$\left\|{h}\right\|_{0}$ $\leq\sum_{i=0}^{k}{\left\|{\frac{f_{i}}{g}}\right\|_{0}}$
$\leq 2\sum_{i=0}^{k}{\left\|{f_{i}}\right\|_{0}\cdot\left(\frac{\deg{f_{i}}}{\deg{g}}\right)^{2}}$
$\leq 2\sum_{i=0}^{k}{\left\|{f_{i}}\right\|_{0}^{3}\cdot\left(\frac{b(f,g)}{\deg{g}}\right)^{2}}$
$\leq 2\left\|{f}\right\|_{0}^{3}\left(\frac{b(f,g)}{\deg{g}}\right)^{2}$
$=\frac{1}{2}\left\|{f}\right\|_{0}^{3}\log^{6}{(3\deg{g})}\left(\log{\left\|{f}\right\|_{\infty}}+6\log{\deg{f}}+10\right)^{2},$
which completes the proof of the theorem. ∎
5 Divisibility Testing over $\mathbb{Z}$
In this section, we obtain a polynomial time algorithm for testing divisibility by binomials and cyclotomic-free trinomials over $\mathbb{Z}$ using polynomial division algorithm (see [CLO13, Chapter 1.5] or [vzGG13, Chapter 2.4]), also called long division or Euclidean division algorithm. Recall that Theorem 1.12 and Theorem 1.13 bound the sparsity of the quotient when the division is exact. Thus, if during the algorithm the sparsity of the quotient exceeds these bounds, we can terminate the algorithm and declare that the division is not exact. Similarly, we use Theorem 1.8 to terminate the algorithm when the coefficients of the quotient become too large. In other words, if the sparsity of the quotient is bounded by $s$, and its coefficients are bounded (in absolute value) by $c$, then we execute $s$ steps of the division algorithm and return True if and only if the remainder is zero. If at any step $\left\|{h}\right\|_{\infty}>c$, we return False.
Algorithm 1 Bounded Polynomial Division over $\mathbb{Z}$
$g,f\in\mathbb{Z}[x]$, a bound $s\in\mathbb{N}$ on $\|f/g\|_{0}$, and a bound $c\in\mathbb{N}$ on $\|f/g\|_{\infty}$.
True and a quotient $f/g$ if $g\mid f$; False otherwise.
  1. 1.
    Set $q=0$, $r=f$, $i=1$, and $t=0$.
  2. 2.
    While $\deg{r}\geq\deg{g}$, $|LC(t)|\leq c$, and $i\leq s$:
    1. (a)
      If $LC(g)\nmid LC(r)$, return False.
    2. (b)
      Set $t=LT(r)/LT(g)$.
    3. (c)
      Set $q=q+t$.
    4. (d)
      Set $r=r-t\cdot g$.
    5. (e)
      Set $i=i+1$.
  3. 3.
    If $r=0$ return True and the quotient $q$; Otherwise, return False.
The correctness of Algorithm 1 is due to the fact that in every step of the loop, we reveal a single term of the quotient. By the correctness of the bounds $c$ and $s$, the coefficient of the term cannot exceed $c$, and the number of terms cannot be larger than $s$, unless $g\nmid f$.
Proposition 5.1.
The time complexity of Algorithm 1 is
$\tilde{O}\left((\left\|{f}\right\|_{0}+s\cdot\left\|{g}\right\|_{0})\cdot(\log{\left\|{g}\right\|_{\infty}}+\log{\left\|{f}\right\|_{\infty}}+\log{c}+\log\deg{f})\right)$
Proof.
Each iteration of the While loop requires $O(\left\|{g}\right\|_{0})$ multiplications, additions and divisions of integers whose absolute values are at most
$B\leq\left\|{f}\right\|_{\infty}+s\cdot c\cdot\left\|{g}\right\|_{\infty}.$
Hence, its bit complexity is at most $b=\lceil\log_{2}B\rceil$. These arithmetic operations can be performed in time
$O(b\log b)=\tilde{O}(\log s+\log c+\log\left\|{g}\right\|_{\infty}+\log\left\|{f}\right\|_{\infty})$
(see [HvdH21]). In addition, we perform operations on the exponent that require $O(\log\deg f)$ time and we have to store both $q$ and $r$ in memory. Storing $q$ requires storing $s$ terms, each term has degree at most $\deg f$ and its coefficient is at most $c$ in absolute value. Thus, the cost of storing $q$ is at most $s(\log_{2}{\deg{f}}+\log_{2}{c})$. We next bound the complexity of storing the remainder, $r$. In each iteration of the While loop the number of terms of $r$ increases (additively) by at most $\left\|{g}\right\|_{0}-1$. Similarly, $\left\|{r}\right\|_{\infty}$ increases (additively) by at most $\leq\left\|{g}\right\|_{\infty}\cdot c$ in each step. Hence, in every step of the algorithm, it holds that
$\left\|{r}\right\|_{0}$ $\leq\left\|{f}\right\|_{0}+s(\left\|{g}\right\|_{0}-1),$
$\log_{2}{\left\|{r}\right\|_{\infty}}$ $\leq\log_{2}{\left\|{f}\right\|_{\infty}}+\log_{2}{\left\|{g}\right\|_{\infty}}+\log_{2}{s}+\log_{2}{c}.$
This implies that the representation size of $r$ is at most
$\text{size}\left(r\right)$ $=\left\|{r}\right\|_{0}\cdot(\log_{2}{\left\|{r}\right\|_{\infty}}+\log_{2}{\deg{r}})$
$\leq(\left\|{f}\right\|_{0}+s\cdot(\left\|{g}\right\|_{0}-1))\cdot(\log_{2}{\left\|{f}\right\|_{\infty}}+\log_{2}{\left\|{g}\right\|_{\infty}}+\log_{2}{s}+\log_{2}{c}+2\log_{2}{\deg{f}}).$
Combining all these estimates we get that the running time is at most
$\tilde{O}\left((\left\|{f}\right\|_{0}+s\cdot\left\|{g}\right\|_{0})\cdot(\log{\left\|{g}\right\|_{\infty}}+\log{\left\|{f}\right\|_{\infty}}+\log{c}+\log\deg{f})\right).\qed$
The proof of Theorem 1.10 easily follows.
Proof of Theorem 1.10.
Let $s=\deg(f)\geq\left\|{f/g}\right\|_{0}$ and $c=\log_{2}{\left\|{f}\right\|_{\infty}}+\tilde{O}(\|g\|_{0})\cdot\log{\deg{f}}\geq\left\|{f/g}\right\|_{\infty}$ be as in Equation (1.9). Note that $\deg f$ is a very rough upper bound on $\|f/g\|_{0}$, but as the division is exact the algorithm will terminate after $\left\|{f/g}\right\|_{0}$ iterations. Thus, we may assume in the upper bound calculations that $s=\left\|{f/g}\right\|_{0}$. Substituting these values to 5.1 (and observing that $\left\|{f}\right\|_{0}\leq\left\|{g}\right\|_{0}\left\|{f/g}\right\|_{0}$), gives the desired result. ∎
Other corollaries to 5.1 concern dividing by a binomial and a cyclotomic-free trinomial.
Corollary 5.2.
Let $g,f\in\mathbb{Z}[x]$ such that $g$ is a binomial. Then, there exists an algorithm that decides whether $g\mid f$ (and if so, computes the quotient $f/g$) in
$\tilde{O}(\left\|{f}\right\|_{0}\cdot\log{\left\|{f}\right\|_{\infty}}\cdot(\log{\left\|{g}\right\|_{\infty}}+\log{\left\|{f}\right\|_{\infty}}+\log{\deg{f}}))$
time. In particular, the runtime complexity of the algorithm is $\tilde{O}(n^{2})$ in the representation size.
Proof.
If $g=c(x^{m}\pm 1)$ for some $m\in\mathbb{N}$ and $c\in\mathbb{Z}$, we can test $c\mid f$ and compute the remainder $f\mod x^{m}\pm 1$ directly, by translating each monomial $x^{k}$ in $f$ to $x^{k\bmod m}$. Otherwise, execute Algorithm 1 with $s$ as in Theorem 4.6, and $c$ as in Theorem 3.10. Since
$s$ $=O(\left\|{f}\right\|_{0}(\log{\left\|{f}\right\|_{0}}+\log{\left\|{f}\right\|_{\infty}}))\;$
$\log{c}$ $=O(\log{\deg{f}}+\log{\left\|{f}\right\|_{\infty}}),$
we conclude the result by 5.1. ∎
Corollary 5.3.
Let $g,f\in\mathbb{Z}[x]$ such that $g$ is a cyclotomic-free trinomial. Then, there exists an algorithm that decides whether $g\mid f$ (and if so, computes the quotient $f/g$) in
$\tilde{O}(\text{size}\left(f\right)^{2}\cdot\log^{6}{\deg{g}}\cdot(\text{size}\left(f\right)+\left\|{f}\right\|_{0}\log{\left\|{g}\right\|_{\infty}}))$
time. In particular, the runtime complexity of the algorithm is $\tilde{O}(n^{10})$ in the representation size.
Proof.
Execute Algorithm 1 with $s$ as in Theorem 4.17, and $c$ as in Theorem 3.10. Since
$s$ $=O(\left\|{f}\right\|_{0}\cdot\text{size}\left(f\right)^{2}\cdot\log^{6}{\deg{g}})$
$\log{c}$ $=O(\log{\deg{f}}+\log{\left\|{f}\right\|_{\infty}}),$
the result follows from 5.1. ∎
Remark 5.4 (Divisibility over $\mathbb{Q}$).
To handle divisibility testing over $\mathbb{Q}$ instead of $\mathbb{Z}$, we multiply $f$ by $LC(g)^{s}$, where $s$ is the number of steps as in Algorithm 1. This ensures all coefficients of the quotient are in $\mathbb{Z}$, and increases $\text{size}\left(f\right)$ by a factor of $LC(g)\cdot\log_{2}{s}$.
Remark 5.5 (Comparison to [GGPdC21]).
When restricted to the trinomial case, Theorem 3.8 of [GGPdC21] implies that there exists a polynomial time algorithm that decides $g\mid f$ when $g$ is a trinomial and $\deg{f}=O(\deg{g})$. Our algorithm works for any degree but requires $g$ to be cyclotomic-free. Without assuming that $g$ is cyclotomic free, our algorithm works as long as $\deg{f}=\tilde{O}(\deg{g})$.
6 Divisibility Testing over Finite Fields
Throughout this section, we assume without loss of generality that $g$ is monic: if $R$ is a field, we can normalize $g$ without changing the fact that $g\mid f$.
6.1 Testing Divisibility by a Binomial
In this subsection, we prove Theorem 1.16 and 1.17. Testing divisibility by a binomial is a relatively easy task: we observe that in this case, $f\bmod g$ must be sparse. Hence, we can calculate the remainder polynomial of the division of $f$ by $g$, and test if it is equivalent to the zero polynomial.
Observation 6.1.
Denote $g=x^{m}-a$. Then,
$x^{e}\bmod g=x^{r}a^{q}$
where $q$ and $r$ are the quotient and the remainder of the integer division of $e$ by $m$, respectively. In other words, $q$ and $r$ are the unique integers such that $e=qm+r$ and $0\leq r<m$.
Since the sparsity of a sum of polynomials is bounded by the sum of the corresponding sparsities, the above observation immediately implies that $\left\|{f\bmod g}\right\|_{0}\leq\left\|{f}\right\|_{0}$ for any $f\in\mathbb{F}[x]$.
We wrap up the divisibility testing algorithm as follows: Let $f=\sum_{i}{c_{i}x^{e_{i}}}$ and $g=x^{m}-a$. For each monomial of $f$, we calculate $c_{i}\cdot x^{e_{i}}\bmod g$ by first calculating $r$ and $q$ using Euclidean division and then calculating $c_{i}\cdot a^{q}$ using repeated squaring. By adding the results, we get an $\tilde{O}(n^{2})$ time algorithm to compute $f\bmod g$ for any binomial $g$, where $n=\text{size}\left(f\right)+\text{size}\left(g\right)$.
Binomials are a special case of a broader class of polynomials in which the exponents have a large common denominator. If $g=x^{m}-a$, then $g=\tilde{g}(x^{m})$ where $\tilde{g}=x-a$. 4.5 implies a divisibility testing algorithm for a more general case: if $g=\ell(x^{m})$ where $\ell$ is a low-degree polynomial, then we can test whether $g\mid f$ in time that is polynomial in $\text{size}\left(f\right)$ and $\deg{\ell}$, by computing all the remainders $u_{j}\bmod\ell$. We note that in the binomial case, when $g=x^{m}-a$, 4.5 combined with 2.2 implies that $g\mid f$ if and only if $u_{j}(a)=0$ for all $0\leq j<m$.
Algorithm 2 Special Case of Polynomial Divisibility Testing over a Finite Field
$f,g\in\mathbb{F}[x]$ such that $g=\ell(x^{m})$ and $\mathbb{F}$ is a finite field.
True if $g\mid f$; False otherwise.
  1. 1.
    Apply the preprocessing step in Section 2.3 to get $\text{ord}_{0}{g}=0$ and that $g$ is monic.
  2. 2.
    Denote $f=\sum_{i}{f_{i}x^{i}}$ and $u_{j}$ as in 4.5.
  3. 3.
    For all $j\in\left\{i\bmod m:f_{i}\neq 0\right\}$:
    1. (a)
      If $u_{j}\bmod\ell\neq 0$ return False.
  4. 4.
    Return True.
The correctness of the algorithm is a direct consequence of 4.5.
Proposition 6.2 (Efficiency).
The time complexity of Algorithm 2 is,
$\tilde{O}(\left\|{f}\right\|_{0}\log{\deg{f}}\cdot\log{|\mathbb{F}|}\cdot\deg{\ell}).$
Proof.
The sum of sparsities of $u_{j}$ for all $0\leq j<m$ is bounded by $\left\|{f}\right\|_{0}$. Computing the remainder $x^{e}\mod\ell$ of a monomial of degree $\leq\deg{f}$ can be done using repeated squaring, which takes $\tilde{O}(\log{\deg{f}}\cdot\log{|\mathbb{F}|}\cdot\deg{\ell})$ time. ∎
We note that 4.5 can be used for polynomials over $\mathbb{Z}$ as well; however, as $u_{j}\bmod\ell$ may have exponentially large coefficients, this does not immediately imply a similar algorithm. To overcome this difficulty, see 4.16 for testing divisibility by a low degree polynomial over $\mathbb{Z}$.
6.2 Testing Divisibility by a Pentanomial
In this section, we prove Theorem 1.18. We first study the case where the degrees of $f$ and $g$ are very close, $\deg{f}<\deg{g}\cdot(1+\frac{1}{\left\|{g}\right\|_{0}-1})$. We proceed to show a reduction from the case of $\deg{f}=\tilde{O}(\deg{g})$ to the previous case.
6.2.1 The Case of $\deg{f}<\deg{g}\cdot(1+\frac{1}{\left\|{g}\right\|_{0}-1})$
We next show that when $\deg{f}<\deg{g}\cdot(1+\frac{1}{\left\|{g}\right\|_{0}-1})$ we can break the polynomials $g$ and $f$ into two polynomials $g_{0},g_{0}$ and $f_{0},f_{1}$, respectively, such that $g_{0}\mid f_{0}$ and $g_{1}\mid f_{1}$. We start with a simple lemma that shows that sparse polynomials have large gaps between consecutive monomials, and then conclude a simple test for divisibility of pentanomials.
Claim 6.3 (Large Gap Exists).
Let $R$ be a UFD. Let $g\in R[x]$ such that $\text{ord}_{0}{g}=0$, and $\left\|{g}\right\|_{0}\geq 2$. Then, there exist $d\in\mathbb{N}$ and nonzero polynomials $g_{0},g_{1}$ such that $g=g_{0}+x^{d}g_{1}$ with $d-\deg{g_{0}}\geq\frac{\deg{g}}{\left\|{g}\right\|_{0}-1}$.
Proof.
Let $e_{1},\dots,e_{\left\|{g}\right\|_{0}}$ be the exponents that correspond to the non-zero coefficients of $g$, and observe that by our assumption $e_{1}=0$ and $e_{\left\|{g}\right\|_{0}}=\deg{g}$. Since
$\sum_{i=2}^{\left\|{g}\right\|_{0}}{(e_{i}-e_{i-1})}=e_{\left\|{g}\right\|_{0}}-e_{1}=\deg{g},$
there exists $2\leq i\leq\left\|{g}\right\|_{0}$ such that,
$e_{i}-e_{i-1}\geq\frac{1}{\left\|{g}\right\|_{0}-1}\sum_{i=2}^{\left\|{g}\right\|_{0}}{(e_{i}-e_{i-1})}=\frac{\deg{g}}{\left\|{g}\right\|_{0}-1}$
Finally, we set $d=e_{i}$ and $g_{0}=f\mod x^{d}$. The claim follows as $\deg{g_{0}}=e_{i-1}$. ∎
Lemma 6.4.
Let $R$ be a UFD, and $d\in\mathbb{N}$. Let $f,g\in R[x]$ such that,
  1. 1.
    $g=g_{0}+x^{d}g_{1}$ with $g_{0},g_{1}\neq 0$ and $\deg{g_{0}}<d$.
  2. 2.
    $f=f_{0}+x^{d}f_{1}$ with $\deg{f_{0}}<d$.
  3. 3.
    $\deg{f}-\deg{g}<d-\deg{g_{0}}$.
Then, $g\mid f$ if and only if $g_{0}f_{1}=g_{1}f_{0}$ and $g_{i}\mid f_{i}$ for some $i\in\left\{0,1\right\}$.
Proof.
Let $0\neq h\in R[x]$ such that $f=gh$. Thus,
$f_{0}+x^{d}f_{1}=f=gh=(g_{0}+x^{d}g_{1})h=g_{0}h+x^{d}g_{1}h,$
and observe that by the assumption,
$\deg{g_{0}h}$ $=\deg{g_{0}}+\deg{h}$
$=\deg{g_{0}}+\deg{f}-\deg{g}$
$<d.$
Thus, by comparing powers of $x$ we conclude that $f_{0}=g_{0}h$ and $f_{1}=g_{1}h$, and in particular $g_{0}\mid f_{0}$ and $g_{1}\mid f_{1}$. By combining the two equations above it holds that $f_{0}g_{1}h=g_{0}hf_{1}$ and thus $g_{0}f_{1}=g_{1}f_{0}$. For the converse, assume without loss of generality that $g_{0}\mid f_{0}$, and let $0\neq h\in R[x]$ such that $f_{0}=g_{0}h$. Since $g_{0}f_{1}=g_{1}f_{0}$ it holds that $g_{0}f_{1}=f_{0}g_{1}=g_{0}hg_{1}$ and so we conclude that $f_{1}=g_{1}h$. Hence
$f=f_{0}+x^{d}f_{1}=g_{0}h+x^{d}g_{1}h=h(g_{0}+x^{d}g_{1})=hg,$
which implies $g\mid f$. ∎
Finally, by combining 6.4 with 6.3 we obtain the following algorithm for the case in which $g$ is a pentanomial and $\deg{f}<\deg{g}\cdot(1+\frac{1}{\left\|{g}\right\|_{0}-1})$.
Algorithm 3 Divisibility Testing by a Pentanomial over a Finite Field (The Case of Low-Degree)
$f,g\in\mathbb{F}[x]$ such that $\left\|{g}\right\|_{0}\leq 5$ and $\deg{f}<\deg{g}\cdot(1+\frac{1}{\left\|{g}\right\|_{0}-1})$.
True if $g\mid f$; False otherwise.
  1. 1.
    Apply the preprocessing step in Section 2.3 to get $\left\|{g}\right\|_{0}\geq 2$ and $\text{ord}_{0}{g}=0$.
  2. 2.
    Find $d$ such that $g=g_{0}+x^{d}g_{1}$ with $g_{0},g_{1}\neq 0$ and
    $d-\deg{g_{0}}\geq\frac{\deg{g}}{\left\|{g}\right\|_{0}-1}.$
  3. 3.
    Let $f=f_{0}+x^{d}f_{1}$ such that $\deg{f_{0}}<d$.
  4. 4.
    If $f_{0}\cdot g_{1}\neq g_{0}\cdot f_{1}$ then return False.
  5. 5.
    Test $g_{i}\mid f_{i}$ for $i\in\left\{0,1\right\}$ such that $\left\|{g_{i}}\right\|_{0}\leq 2$, using Algorithm 2.
Proposition 6.5 (Correctness).
Algorithm 3 returns True if and only if $g\mid f$
Proof.
The existence of $d$ in item 2 is a corollary of the assumptions and 6.3. The existence of $g_{i}$ in item 5 with $\left\|{g_{i}}\right\|\leq 2$ is due to the fact that $\left\|{g_{0}}\right\|_{0}+\left\|{g_{1}}\right\|_{0}=\left\|{g}\right\|_{0}\leq 5$. The algorithm correctness follows from 6.4 since $\deg{f}-\deg{g}<\frac{\deg{g}}{\left\|{g}\right\|-1}\leq d-\deg{g_{0}}$. ∎
Proposition 6.6 (Efficiency).
The time complexity of Algorithm 3 is,
$O(\left\|{f}\right\|_{0}\log{\deg{f}}\log{|\mathbb{F}|}).$
In particular, it runs in $O(n)$ for constant-size fields ($O(n^{2})$ in the representation size).
Proof.
Steps 2-4 can be computed by iterating over the coefficients and exponents of $g$ and $f$. This takes $O(\left\|{f}\right\|_{0}(\log{\deg{f}}+\log{\left|\mathbb{F}\right|}))$ to execute, due to the assumption that $\left\|{g_{0}}\right\|_{0},\left\|{g_{1}}\right\|_{0}\leq 5$. Step 5 is computed using Algorithm 2 (with $\deg{\ell}=1$ as the chosen $g_{i}$ is a binomial), which takes $O(\left\|{f}\right\|_{0}\log{\deg{f}}\cdot\log{\left|\mathbb{F}\right|})$ time to execute. ∎
6.2.2 The Case of $\deg{f}\leq\tilde{O}(\deg{g})$
In this section, we extend the results of the previous section by showing a reduction to the case $\deg{f}<\deg{g}\cdot(1+\frac{1}{\left\|{g}\right\|_{0}-1})$ that was handled in Section 6.2.1.
Lemma 6.7 (Degree Reduction).
Let $p$ be a prime, and let $k\in\mathbb{N}$. Let $f,g\in\mathbb{F}_{p}[x]$ such that $\text{ord}_{0}{g}=0$ and define:
$v$ $=f\cdot g^{p^{k}-1}\;(=\sum_{i}v_{i}x^{i}),$
$u_{j}$ $=\sum_{i\;\equiv_{p^{k}}\;j}{v_{i}x^{(i-j)/p^{k}}}.$
Then, $g\mid f$ if and only if $g\mid u_{j}$ for all $0\leq j<p^{k}$.
Proof.
By 2.1 and 2.4 it holds that
$g\mid f\iff g^{p^{k}}\mid f\cdot g^{p^{k}-1}\iff g(x^{p^{k}})\mid v.$
The claim follows from 4.5. ∎
6.7 gives rise to a natural algorithm that reduces divisibility testing into multiple divisibility tests of polynomials of lower degree. However, this also increases the sparsity of the resulting polynomials. Thus, we need to bound the total sparsity of the polynomials $\left\{u_{j}:0\leq j<p^{k}\right\}$ and $g^{p^{k}-1}$.
Lemma 6.8.
Let $p$ be a prime, and let $k\in\mathbb{N}$. Let $g\in\mathbb{F}_{p}[x]$. Then,
$\left\|{g^{p^{k}-1}}\right\|_{0}\leq p^{k(\left\|{g}\right\|_{0}-1)}.$
Proof.
The polynomial $g^{p^{k}-1}$ can be presented as a product of $k$ polynomials,
$g^{p^{k}-1}=\frac{g^{p^{k}}}{g}=\prod_{j\in[k]}{\frac{g^{p^{j}}}{g^{p^{j-1}}}}.$
By 2.4 it holds that,
$\frac{g^{p^{j}}}{g^{p^{j-1}}}=g^{p^{j}-p^{j-1}}=g^{p^{j-1}(p-1)}=g(x^{p^{j-1}})^{p-1},$
and in particular, $\left\|{\frac{g^{p^{j}}}{g^{p^{j-1}}}}\right\|_{0}=\left\|{g^{p-1}}\right\|_{0}$. Thus, we obtain the following bound on the sparsity of the polynomial $g^{p^{k}-1}$:
$\left\|{g^{p^{k}-1}}\right\|_{0}\leq\prod_{j\in[k]}{\left\|{\frac{g^{p^{j}}}{g^{p^{j-1}}}}\right\|_{0}}=\prod_{j\in[k]}{\left\|{g^{p-1}}\right\|_{0}}=\left\|{g^{p-1}}\right\|_{0}^{k}.$
The result follows from a bound on the sparsity of $g^{p-1}$. It holds that,
$\left\|{g^{p-1}}\right\|_{0}\leq\binom{\left\|{g}\right\|_{0}+p-2}{\left\|{g}\right\|_{0}-1}\leq p^{\left\|{g}\right\|_{0}-1}$
where the first inequality is due to 2.3 and the second is due to 2.10. ∎
Proposition 6.9.
Let $p$ be a prime and let $k\in\mathbb{N}$. Let $f,g\in\mathbb{F}_{p}[x]$ such that $\text{ord}_{0}{g}=0$ and let $\left\{u_{j}:0\leq j<p^{k}\right\}$ defined as in 6.7. Then,
  1. 1.
    $\deg{u_{j}}\leq\deg{g}+\frac{\deg{f}-\deg{g}}{p^{k}}$ for all $0\leq j<p^{k}$.
  2. 2.
    $\sum_{0\leq j<p^{k}}{\left\|{u_{j}}\right\|_{0}}\leq\left\|{f}\right\|_{0}\cdot p^{k(\left\|{g}\right\|_{0}-1)}$.
Proof.
Let $v$ be defined as in 6.7. It holds that
$f\cdot g^{p^{k}-1}=v=\sum_{0\leq j<p^{k}}{x^{j}u_{j}(x^{p^{k}})}.$ (6.10)
Let $0\leq j<p^{k}$. We bound the degree of $u_{j}$ by first observing the degree of $v$ is
$\deg{v}=\deg{(f\cdot g^{p^{k}-1})}=\deg{f}+\deg{g}\cdot(p^{k}-1).$
Thus,
$\deg{u_{j}}$ $\leq\frac{\deg{v}}{p^{k}}$
$=\frac{\deg{f}+\deg{g}\cdot(p^{k}-1)}{p^{k}}$
$=\deg{g}+\frac{\deg{f}-\deg{g}}{p^{k}}$
which proves the first item. By Equation 6.10 we conclude, together with 6.8,
$\sum_{0\leq j<p^{k}}{\left\|{u_{j}}\right\|_{0}}$ $=\left\|{v}\right\|_{0}$
$\leq\left\|{f}\right\|_{0}\cdot\left\|{g^{p^{k}-1}}\right\|_{0}$
$\leq\left\|{f}\right\|_{0}\cdot p^{k(\left\|{g}\right\|_{0}-1)}$
which concludes the proof. ∎
Finally, 6.7 together with 6.9 give rise to the following algorithm.
Algorithm 4 Divisibility Testing by Pentanomial over a Finite Field (The Case of Medium-Degree)
$f,g\in\mathbb{F}_{p}[x]$ such that $p$ is a prime, $\left\|{g}\right\|_{0}\leq 5$, and $\deg{f}=\tilde{O}(\deg{g})$.
True if $g\mid f$; False otherwise.
  1. 1.
    Apply the preprocessing step in Section 2.3 to get $\text{ord}_{0}{g}=0$.
  2. 2.
    Compute $k=\lceil\log_{p}{\left\|{g}\right\|_{0}}+\log_{p}{(\frac{\deg{f}}{\deg{g}}-1)}\rceil$.
  3. 3.
    Compute $v=fg^{p^{k}-1}=\sum_{i}v_{i}x^{i}$.
  4. 4.
    For all $j\in\left\{i\bmod p^{k}:v_{i}\neq 0\right\}$:
    1. (a)
      Let
      $u_{j}=\sum_{i\;\equiv_{p^{k}}\;j}{v_{i}x^{(i-j)/p^{k}}}$
      (as in 6.7).
    2. (b)
      Test $g\mid u_{j}$ using Algorithm 3 and return False if fails.
  5. 5.
    Return True.
Proposition 6.11 (Correctness).
Algorithm 4 returns True if and only if $g\mid f$.
Proof.
Observe that $p^{k}\geq\left\|{g}\right\|_{0}(\frac{\deg{f}}{\deg{g}}-1)$. By 6.9,
$\deg{u_{j}}$ $\leq\deg{g}+\frac{\deg{f}-\deg{g}}{p^{k}}$
$\leq\deg{g}+\frac{\deg{f}-\deg{g}}{\left\|{g}\right\|_{0}(\frac{\deg{f}}{\deg{g}}-1)}$
$=\deg{g}+\frac{\deg{g}}{\left\|{g}\right\|_{0}},$
which means that the assumptions of Algorithm 3 hold. Thus, the correctness of the algorithm follows from 6.7 and the correctness of Algorithm 3. ∎
Proposition 6.12 (Efficiency).
The time complexity of Algorithm 4 is,
$O\left(\left\|{f}\right\|_{0}\left(p\left(\frac{\deg{f}}{\deg{g}}-1\right)\right)^{\left\|{g}\right\|_{0}-1}\log{\deg{g}}\log{p}\right).$
In particular, it runs in $poly(n)$ time.
Proof.
By the correctness of the algorithm, it holds that
$\deg{u_{j}}\leq\deg{g}+\frac{\deg{g}}{\left\|{g}\right\|_{0}}\leq 2\deg{g}.$
Observe that,
$p^{k}\leq p\left\|{g}\right\|_{0}\left(\frac{\deg{f}}{\deg{g}}-1\right).$
Hence, by 6.9,
$\sum_{0\leq j<p^{k}}{\left\|{u_{j}}\right\|_{0}}$ $\leq\left\|{f}\right\|_{0}\cdot p^{k(\left\|{g}\right\|_{0}-1)}$
$\leq\left\|{f}\right\|_{0}\left(p\left\|{g}\right\|_{0}\left(\frac{\deg{f}}{\deg{g}}-1\right)\right)^{\left\|{g}\right\|_{0}-1}.$
Thus, the total sparsity of all of the polynomials $u_{j}$ in the invocations to Algorithm 3 is bounded by the above expression. The total time complexity follows from the time complexity of Algorithm 3. ∎
Acknowledgment
The authors would like to thank Bruno Grenet for his helpful comments and for suggesting 1.11. We would also like to thank the anonymous referees for their comments.